A Complete Bibliography of the Journal of Human–Robot Interaction (JHRI) and ACM

Transactions on Human-Robot Interaction (THRI)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

30 August 2023 Version 1.04

Title word cross-reference

19 [FSHR^{+}21].

2014 [SPGB21].

Ability [RSB22, YZL^{+}16]. Abject [BKH23]. Acceptability [GAG^{+}21]. Acceptance [MTCKS21]. Accepting [SV22]. achieve [HM13]. ACM [JS18, JS19, SJ18, SJ22]. Acquisition [FSLN19, LMD14]. ACT [THH^{+}13]. ACT-R [THH^{+}13]. ACT-R/E [THH^{+}13]. Action [WE20]. Activation [BBM^{+}21]. Active [LKL23, FSP15]. Activities [MNM18]. activity [CBO^{+}16, JBF^{+}14]. acts [BWS17]. Adapt [SHSK20, MM16]. Adaptation [GRSA23, VPR^{+}22]. Adapting [CBK^{+}18]. adaptive [HLWV14]. Admonishments [ESK23]. Adults [BFW^{+}20, BBM^{+}21, ECL^{+}22, GAG^{+}21, GAL^{+}22, JL19, PPRR20, MN13]. advances [Goo13]. Aesthetic [Lju23]. Aesthetics [Jør23]. Affect [SGJ^{+}22]. Affect-Aware [SGJ^{+}22]. Affective [BMB23b, HJ23, PHZ23, LMD14]. Affects [BJ23, WAAR20]. against [SSQS21]. Age [CCL^{+}19]. agency [JL16, SDFSM17]. Agent [RTSL23]. Agents [LT23, WTC^{+}21]. Aging [LR18, LR23]. AI [MK22, Vel18]. AI-driven [MK22]. Aid [MNM18]. aided [Tan15]. Aids [JL19]. air [DLDS^{+}15]. Algorithm [HHHA19, KF18]. Algorithms [FN22]. Allocation [JDS^{+}21]. Alternatives [LCTRS22]. Alzheimer [PPRR20]. Among [LT23]. analog [PSB^{+}16]. Analysing [BEBT22]. Analysis [ASJD19, SST21, CT12, DWM22]. Android [Har16, KTTO13]. animal [PSB^{+}16]. Animation [STH19]. Anthropomorphic [OH22, SS18, SH21]. Anthropomorphism [CK20, RMO23, YZL^{+}16]. Anthropomorphized [SH21]. Anthropomorphizing [Fis22]. Appearance [AGND23, MHI16]. Appearances [JLV^{+}21]. Application [ZFS^{+}20, DSPM15]. Applications [BOMLT23, MAC15]. Applying [Lju23]. Approach [JLV^{+}21, SZA22, Aug14, BRSTV17, BFL^{+}17, DSMS17]. Appropriate [ESK23]. Arab [KAG21]. Architecture [PG22, THH^{+}13]. aren’t [WM16]. Argumentation [SA15]. Argumentation-based [SA15]. Arms [BSH^{+}23]. arousing [LJR17]. Art [MK22]. Artificial [HCR18, MK22]. Artists [MK22]. Asimo [JL16]. Asking [HJL22]. aspects [You17]. Assembly [MTCKS21]. assessing [GCMC15]. Assessment [HHHA19, KBvdP23, ME20, NAC^{+}22]. assist [MBK^{+}12]. Assistance [CTA21, ITI15, NWS15, SD15]. Assisted [JDS^{+}21, Lju23, ZCG^{+}20, Ran20]. Assistive [BOMLT23, CTA21, GAL^{+}22, JA20, LR18, LPPOG23, MNM18, SGJ^{+}22, SOG^{+}22, WTDS20, FM13, HLWV14, MN13, SDFSM17, TOW^{+}13]. At-Home [BBM^{+}21]. Attachment [LCTRS22]. Attention [OH22, SSK^{+}19, RFLM13]. attentive [KF21]. Attitude [CCL^{+}19, WML15]. attitude-aware [WML15]. Attitudes [BOMLT23, TOW^{+}13]. attributing [JL16]. Attribution [TdGZ22]. Attributions [KBvdP23]. Augmented [CHS^{+}22, WPC^{+}23]. Autism [JBJ^{+}20, SGJ^{+}22, KPSS12, SDFSM17]. Automated [SZA22]. Automatic [NMW^{+}21]. Automation [OPK^{+}22, Asa16]. Autonomous [BSH^{+}23, WTDS20, RFLM13, WM16]. Autonomy [CHS21, FN22, JA20, KNF^{+}22, ZZKC19, BFR14, MBK^{+}12]. Avatars [SV22]. Average [RC15]. Average-position [RC15]. avoidance [ITI15]. Avoiding [BKH23]. Aware [CNS^{+}20, SGJ^{+}22, SZA22, WML15]. awareness [PCB^{+}17, PJ17].

Back [RPB21]. Back-off [RPB21].

Balancing [Mat18]. Based

[FM22, KDC^{+}21, WKP^{+}23, BWS17, BTLY15, DLG^{+}19, DKI19, FSM12, HBAZ15, KKC^{+}22, MHI16, NWS15, RSB22, SA15, WS13]. bases [DSPM15]. Be [HH23, MK22]. beginnings [RTSL23]. Behavior [DKI19, Fis22, HGA^{+}21, HY23, JDS^{+}21, KLE19, KF21, LKL23, ORPW19, HM13, MM16]. Behavioral [BK22, BBM^{+}21, LZ17]. Behaviors [MNM18]. behaviour [DSMS17]. Belief [TZ21]. Bender [dSCV^{+}13]. Benefits [LCTRS22]. Best [BHH^{+}23, SGHb^{+}23]. Better [CTA21, FS17, FGCB23, Sch14]. between [NWS15, SDFSM17]. Beyond [KLE19, NACB20]. Bias [HB18, DSMS17]. Blossom [SH19]. Bodily [Lju23]. Body [LPPOG23, PHZ23, SS18, LJR17]. Bonding [MvdPD23]. bonds [BBR^{+}13]. Boosts [VPR^{+}22]. Both [BHH^{+}23]. Brain [ME20]. Brave [Bro18]. break [SRK14]. Breed [WRPZ21]. Bridging [ASJD19, KPSS12]. Brief [Dau18]. Building [HGA^{+}21, BBR^{+}13]. BUSA [BFL^{+}17].

Call [Jør23, WE20]. calling [SPGB21].

Calls [SVD21]. Can

[FGCB23, HH23, MK22, SHSK20]. Canvases [AORK22]. capabilities [dSCV^{+}13]. Capability [KBvdP23]. capable [MBK^{+}12]. Capacities [vSPKB20]. Care [GHD21, MN13, TM12]. care-receiving [TM12]. Caregiver [OPK^{+}22]. categories [Wag15]. Catharsis [LSB^{+}20]. Causal [HY23]. Cause [SSQS21]. centered [MBK^{+}12, WML15]. Centric [TCR20, KA13, KTTO13]. Challenge [RF23, TZ21]. Challenges [LCTRS22, MBW^{+}23, WTDS20]. Changing [DBK21, HH23]. character [Sch14]. Characterizing [PES20, Mil13]. Charismatic [FNJB20]. Charting [SVKF23]. Chief [SJ18]. Child [LNH23, vSPKB20, BBR^{+}13, CBO^{+}16]. Child-Robot [LNH23]. Children [JBJ^{+}20, SHSK20, SGJ^{+}22, TM12, FSM12, SDFSM17]. Clarification [JW22]. Clarifying [DTT^{+}13]. Classification [BSC18, HS15]. classroom [TM12]. Clinicians [JL19]. Closed [BA18]. Closed-Loop [BA18]. Closing [FM22]. Cloud [KKC^{+}22]. Cloud-based [KKC^{+}22]. Co [AORK22, LNH23]. Co-creation [LNH23]. Co-Design [AORK22]. Co-regulation [LNH23]. coach [FM13]. Coactive [JBF^{+}14]. CoBot [Vel18]. Cognitive [DSPM15, GAG^{+}21, HMGC13, LHP^{+}13, SH21, THH^{+}13]. Collaborating [CHS^{+}22]. Collaboration [CNS^{+}20, FBGL^{+}22, FDC^{+}22, JDS^{+}21, LBA^{+}21, MTCKS21, NKFS18, RTC^{+}23]. Collaborative [BJ23, CRA21, GRSA23, HHHA19, MBZA20, GCMC15]. Collective [CRA21, KF21, RCL^{+}22]. Collision [BA18, ITI15]. Collision-Free [BA18]. Color [LPPOG23]. Colorful [WRPZ21]. Comanipulative [RSB22]. Combining [BHH^{+}23]. Commands [SOCAG^{+}19, DTT^{+}13]. Comment [SVD21]. Commentary [Fis21]. Commitment [VPR^{+}22]. Common [MvdPD23]. Communicating [HY23, LCM^{+}19]. Communication [CCT^{+}18, FES22, NKFS18, NAC^{+}22, SST21, US23, WKP^{+}23, MF15, WBS15]. Communications [BMB^{+}23a]. Communities [KAG21]. Company [CK20]. comparative [ZCW^{+}16]. Comparing [WKP^{+}23, BTLY15]. Competent [MAOTK22]. Competitions [MWIN22]. Complexity [SVD21, WdHVvR23]. Compliance [HFK^{+}22]. Computational [WdHVvR23, WYJ23, FSM12]. Computing [MWJO19]. concept [MBK^{+}12, WM16]. Conceptions [TdGZ22]. Conceptual [KNF^{+}22]. Concerning [SGHb^{+}23]. concerns [WM16]. conditions [PJ17]. Conducting [Fis21, HZ21, SVD21, TKC14]. Conference [SCK^{+}22]. confidence [WYJ23]. configuration [WEG^{+}13]. Confirmed [Fis21]. Conflicts [MHV^{+}21, MPCV13]. Conformity [SSQS21]. connectivity [PCB^{+}17]. Consent [SAS19]. considering [You17]. constraints [Tan15]. Construction [JDS^{+}21, WdHVvR23]. contact [LJR17]. Context [HFK^{+}22]. continue [GCSY16]. Control [FM22, GPFA20, WdHVvR23, ZFS^{+}20, ZZKC19, CT12, DLDS^{+}15, GCMC15, Mil13, MAC15, NWS15, SA15, SD15]. controller [WML15]. controllers [SD15]. Controlling [CT13]. Conversational [PPRR20]. Conversations [ECL^{+}22]. Cooperation [ME20, MF15, RC15]. cooperative [NWS15]. Coordinating [LNG22]. Coordination [GPFA20, MWR22, RC15]. Core [MBW^{+}23]. Corners [HJ23]. Correct [LO20]. Couchsurfing [MvdPD23]. Coupled [WYJ23]. Coupling [LKL23]. course [You17]. Cover [LPPOG23]. Covert [WBS15]. COVID [FSHR^{+}21]. COVID-19 [FSHR^{+}21]. creating [Wag15]. creation [LNH23]. Creative [Bro18]. creatures [OP14]. Critical [LR23, PHZ23]. Crowdsourcing [BDO^{+}13, TKC14]. Cues [SH21, US23, WTC^{+}21]. Cultural [KAG21]. Curiosity [CCL^{+}19, DLG^{+}19].

Curiosity-based [DLG^{+}19]. Current [LLT21]. customizable [DSL13].

Daily [MNM18]. dangerous [Har16]. DARPA [RF23]. Data [DKI19, DBK21, GCSY16]. Data-Driven [DBK21]. Dead [Fis21]. Decision [CNS^{+}20, RSB22]. Decision-making [RSB22]. Deep [MLF^{+}21]. Deictic [BHH^{+}23]. Delivered [GAG^{+}21]. Dementia [GHD21]. demonstration [SD15]. Dependent [DWM22]. Depression [BBM^{+}21]. Description [VWKP22]. Descriptions [RVC22, WSG^{+}22]. Design [AORK22, CBK^{+}18, CHS^{+}22, DRH^{+}23, FGCB23, GAL^{+}22, GRSA23, HJ23, LKL22, LKL23, MIFP22, MWIN22, MAOTK22, MPCV13, MHV^{+}21, MD21, OH22, Pat23, RSB22, WPC^{+}23, dWVK23, Aug14, DSPM15, FLA^{+}13, HF14, JBF^{+}14, MBK^{+}12, PSB^{+}16, RFLM13, WML15]. Designerly [Lju23]. Designing [BSH^{+}23, HJ14, HLWV14, KLE19, LR23, PHZ23, SRK14, WRPZ21, WTDS20, JBF^{+}14, WS13, HH23]. Designs [HFK^{+}22, MBZA20]. Destruction [LSB^{+}20]. determinism [HLWV14]. Development [GAL^{+}22, KNF^{+}22, LMD14, PB18]. Diagnostic [HHHA19]. dialog [DTT^{+}13]. dialogic [MBLD17]. dialogue [SA15]. Dialogues [SSK^{+}19]. Did [DLG^{+}19, SRHE21]. Difference [CCL^{+}19]. Differences [BK22]. Different [BHH^{+}23, ZZKC19]. Dig [BFL^{+}17]. Dimensions [Oka18]. direct [GCMC15]. Directions [LLT21]. Disclosure [NKT20]. Disease [BOMLT23, PPRR20, WTDS20]. dishonest [Har16]. Disorder [JBJ^{+}20]. Display [CHS^{+}22]. dissonance [LHP^{+}13]. Distance [FSM12]. Distance-based [FSM12]. Distributed [WdHVvR23, RC15]. Diverse [SJ19b, MBLD17]. Diversity [FN22]. Do [AGND23, FGCB23, JL19, PES20]. does [JL16]. Domestic [AGND23]. Dominance [LCM^{+}19]. don’t [Asa16]. drama [SCG17]. Driven [DBK21, MK22]. Driver [SZA22, NWS15]. Drone [ZZKC19]. drones [WM16]. During [JW21, FDC^{+}22, MPCV13, MU15]. Dyadic [SCK^{+}22]. Dynamic [LNG22, MBK^{+}12]. dynamic-autonomy [MBK^{+}12]. Dynamically [WSG^{+}22]. Dynamics [JDS^{+}21, WYJ23, CT13].

Ease [CCL^{+}19]. Editorial

[AJS19, JS17, JS19, SJ19a, SJ19b]. Editors [SJ18]. Editors-in-Chief [SJ18]. Education [SBB17, MBLD17, ZCV^{+}17]. EEG [BK22, MHI16]. EEG-based [MHI16]. Effect [Fis21, KF18, LLBN22, LRS18, LPPOG23]. Effective [MBZA20]. Effectiveness [RMO23, WSG^{+}22]. Effects [BSC18, JAG^{+}23, dWVK23, vSPKB20, MHI16, Sch14, YZL^{+}16]. Efficacy [PB18]. Efficient [LKL22]. Effort [FDC^{+}22]. Effortful [VPR^{+}22]. Egocentric [KDC^{+}21]. Elderly [NKT20, FM13, MBK^{+}12]. Element [WPC^{+}23]. Elicit [SOG^{+}22]. Embodied [BAC^{+}23, PHZ23, WTC^{+}21, THH^{+}13]. Embodiment [NACB20, RMO23]. Embodiments [DRH^{+}23]. Emergence [RTSL23]. emergency [ITI15]. Emerging [LLT21]. Emotion [HFK^{+}22]. Emotional [LSB^{+}20, PT20]. Empathic [AOSM^{+}19, OP14]. Empathy [JAG^{+}23, SOG^{+}22]. Enabling [BWS17, JW22, CXJS13, DSPM15, GSF^{+}13, HM13]. Encoded [FM22]. Encouraging [NKT20]. endpoint [MU15]. Engage [MLF^{+}21, SHSK20, CBO^{+}16]. engaged [ECL^{+}22]. engagement [RFLM13]. Engaging [SOCAG^{+}19]. Engineering [FGCB23, Sch14]. Enhancement [ZFS^{+}20]. Enhancements [FSP15]. enhancing [Tan15]. environment [BDO^{+}13, WS13]. environment-based [WS13]. Equally [KBvdP23]. Ergodic [FM22]. Ergonomic [LKL22]. Errors [AGND23, TO21, WAAR20]. Ethnography [CK20]. Evaluate [MWIN22]. Evaluated [KBvdP23]. Evaluating [BWS^{+}22, FN22, JW21, BTLY15]. Evaluation [CHS^{+}22, GAG^{+}21, GRSA23, LKL22, MIFP22, MAOTK22, RPB21, VK19, ZFS^{+}20, BDO^{+}13, RFLM13, SRK14]. Evaluations [AGND23, MBZ^{+}22]. Event [Gär19]. evocative [LZ17]. Evolution [LKL22]. Examining [LCTRS22]. Exceptions [SAS19]. exchange [HS15]. exchanges [PJ17]. Execution [BA18, VWKP22]. exercise [FM13]. Exoskeleton [JL19]. Expectations [DRH^{+}23, RVC22]. Experience [KF18]. Experiences [JBJ^{+}20, Lju23, NKT20]. experimental [KSNL13]. Experiments [Fis21, HZ21, SVD21, TM12]. Explain [HPY21, HY23, SHSK20]. Explainability [TZ21]. Explainable [DDMZ21, WTC^{+}21]. Explaining [AKS21, WDR^{+}20]. Explanation [AKS21, HGA^{+}21]. Explanations [HPY21]. Exploration [JS21, GCSY16]. Exploratory [RKD^{+}21]. Explore [MWR22]. Exploring [SAS19, AGND23]. exposing [You17]. Expressing [LBA^{+}21]. Expressions [PT20]. Expressive [BAC^{+}23, SOCAG^{+}19, VK19, WRPZ21]. Extended [GRSA23]. eye [AS17].

FABRIC [GRSA23]. Facial [PT20]. Facilitate [JBJ^{+}20, LNH23, RPB21]. Facilitating [CCT^{+}18, FSM12]. facility [MN13]. Factors [SZA22]. Factory [BMB^{+}23a]. Fascinating [Vel18]. Features [KF21, KBvdP23]. Feedback [CCT^{+}18, BTLY15]. Feeding [Lju23]. feel [JL16]. Fiction [MWJO19]. Field [AOSM^{+}19, BFW^{+}20, ESK23, FES22, MLF^{+}21, TM12, TOW^{+}13]. Fielded [RF23]. Fighting [SLTK22]. Findings [RKD^{+}21, TdGZ22]. First [SVKF23]. First-Hand [SVKF23]. floor [CT13]. Foe [MTCKS21]. Folk [MK22]. foraging [KSNL13]. force [Asa16, HS15, RFL^{+}15]. Forceful [FDC^{+}22]. forestry [WS13]. Formalism [VWKP22]. Format [VWKP22]. Formation [vSPKB20]. Forms [BHH^{+}23]. Foster [PG22]. fostering [PSB^{+}16]. Foundation [HGA^{+}21, SJ19b]. Four [GAL^{+}22]. Four-year [GAL^{+}22]. Framework [ASJD19, AORK22, BWS^{+}22, GRSA23, JBJ^{+}20, MWR22, MAOTK22, BFR14, TKC14]. Frameworks [Mil13]. Framing [Lju23, WAAR20]. Free [BA18, ZCW^{+}16]. Friend [MTCKS21]. Friendly [ESK23]. Fully [GPFA20, WTDS20]. Functional [CS13]. Functionality [WAAR20]. Future [Dau18, HFK^{+}22, SJ19b, PSB^{+}16, WS13].

G [SVD21]. Game [MD21]. Games [GAG^{+}21, MD21, SA15]. gap [KPSS12]. gaps [SPGB21]. Gaze [SH21, AS17]. GEDIS [ZFS^{+}20]. GEDIS-UAV [ZFS^{+}20]. General [CBK^{+}18, dSCV^{+}13]. General-Purpose [CBK^{+}18, dSCV^{+}13]. Generate [Gär19]. Generated [NSHK23]. Generating [KF21]. Generation [FN22, HGA^{+}21, VK19, Tan15]. German [KAG21]. Gesture [CCT^{+}18, KTTO13]. gesture-centric [KTTO13]. Gestures [BAC^{+}23, BSH^{+}23, BHH^{+}23, MHV^{+}21, dWVK23, MPCV13, ZCW^{+}16]. Glanceable [KF21]. Global [BA18]. globe [GE13]. goals [HM13]. Going [PES20]. Got [BEBT22]. graduate [You17]. Granular [JS21]. Greetings [RTSL23]. ground [BTLY15, WML15]. ground-based [BTLY15]. Grounded [BWS^{+}22, WKP^{+}23]. Groundwork [GPFA20]. Group [AOSM^{+}19, WE20]. Groups [JDS^{+}21, TCR20]. Guidance [ESK23].

Guided [FGT18]. guidelines [Rie12].

H [WLTO23]. Hacking [HB18]. Hand [SVKF23, Tan15, ZCW^{+}16]. Handcrafted [SH19]. handovers [SLD^{+}13]. handrest [FSP15]. Hands [SVKF23, Asa16]. Happens [JW21]. Happy [Bro18]. Haptic [CCT^{+}18, NWS15, Oka18, SFEK15, BTLY15, ITI15, RC15, SD15]. haptics [MF15]. Head [CHS^{+}22, JAG^{+}23]. Head-mounted [CHS^{+}22]. Health [GAL^{+}22]. Healthcare [BOMLT23, JAG^{+}23]. helicopter [DSPM15]. Here [HJL22]. Hesitation [MHV^{+}21, MPCV13]. Heterogeneous [RF23]. historical [OP14]. Hoffman [SVD21, Fis21]. Home [BBM^{+}21, WRPZ21, MBK^{+}12]. homes [TOW^{+}13]. Households [BFW^{+}20]. HRI [BFL^{+}17, PB18, AJS19, Asa16, ASJD19, BEBT22, Bro18, Fis21, GE13, Goo13, KB13, KPSS12, KA13, MF15, MWIN22, MBZA20, Mat18, MBLD17, NMW^{+}21, Rie12, RMO23, RFLM13, SBB17, SJ19b, SAS19, SVD21, VWKP22, WAAR20, You17, ZCV^{+}17]. Hug [BSH^{+}23]. Hugging [BSH^{+}23]. Human [BK22, BSC18, CHS^{+}22, CCT^{+}18, CNS^{+}20, CRA21, DC16, Dau18, DKI19, FSHR^{+}21, FBGL^{+}22, FDC^{+}22, Fis22, FGT18, FN22, FES22, GRSA23, HCR18, HHHA19, HFK^{+}22, HZ21, HF14, HB18, HH23, JA20, JW21, JH18, KG18, KSNL13, KNF^{+}22, KF18, LLBN22, LBA^{+}21, LKL23, LNG22, LO20, LSB^{+}20, MIFP22, MF15, MBZ^{+}22, MWR22, MTCKS21, ME20, MvdPD23, MHV^{+}21, NKFS18, NAC^{+}22, ORPW19, Oka18, OH22, PG22, PT20, PSB^{+}16, PB18, RPB21, RF23, RTC^{+}23, RSB22, RVC22, RCL^{+}22, SJ18, SJ22, SCK^{+}22, SGHb^{+}23, STH19, SVD21, SV22, SZA22, SSK^{+}19, TCR20, TO21, US23, Vel18, WPC^{+}23, WSG^{+}22, WYJ23, WLTO23, vSPKB20, AS17, BFR14, BFL^{+}17, BDO^{+}13, BTLY15, BGJK16, CS13, CXJS13, DTT^{+}13, GSF^{+}13, GCMC15, HBAZ15, HMGC13, HS15, KTTO13, LMD14, LHP^{+}13, MHI16, MPCV13, MU15]. human [PJ17, dSCV^{+}13, SA15, SLD^{+}13, TGS12, Tan15, TKC14, THH^{+}13, Wag15, WBS15, vdBBD^{+}16]. Human-animal [PSB^{+}16]. Human-Autonomy [KNF^{+}22]. Human-Collective [CRA21, RCL^{+}22]. Human-Guided [FGT18]. Human-Human [ORPW19]. Human-like [SV22]. Human-Mobile [CCT^{+}18]. Human-Multiagent [BK22]. Human-Robot [FBGL^{+}22, FDC^{+}22, Fis22, FN22, HH23, LLBN22, LBA^{+}21, LNG22, MIFP22, MBZ^{+}22, MWR22, MvdPD23, MHV^{+}21, OH22, PG22, RPB21, RF23, RTC^{+}23, SJ22, SCK^{+}22, US23, WLTO23, NAC^{+}22, SGHb^{+}23, WPC^{+}23]. Humane [SS18]. Humanoid [SS18, VPR^{+}22, DSMS17, FLA^{+}13, LJR17]. Humans [SST21].

Ideas [FS17]. Identification

[ECL^{+}22, IOLB22]. Identified [BOMLT23]. Identifying [SPGB21]. If [JL16]. Ignoring [SLTK22]. Imitation [DBK21]. Impact [AGND23, FSHR^{+}21, JDS^{+}21, KBvdP23, MvdPD23, OH22, RVC22, RTSL23, SH21, ZZKC19, DSMS17, JS17, MPCV13]. Implementation [KKC^{+}22]. implications [JL16, WBS15]. Implicit [CCL^{+}19]. importance [You17]. Impressions [SVKF23]. improve [MM16, vdBBD^{+}16]. improved [PCB^{+}17]. Improving [HS15, SGHb^{+}23]. In-the-Wild [MvdPD23, RKD^{+}21]. Inanimate [RTSL23]. inaugural [TGS12]. inclusion [JS17]. Increasingly [Vel18]. Incremental [WSG^{+}22]. Independent [DWM22]. indirect [BWS17]. individuals [KPSS12]. Indoor [NMW^{+}21]. Industrial [OH22]. Influence [HFK^{+}22, RCL^{+}22, WE20]. influencing [PSB^{+}16]. Information [DBK21, FM22, HY23, DTT^{+}13, PJ17]. information-theoretic [DTT^{+}13]. Informational [SSQS21]. infrared [CS13]. Initial [SST21]. Initiated [JW21]. initiative [CHS21, DLDS^{+}15]. initiatives [ZCV^{+}17]. Integrating [BMB^{+}23a]. Intelligence [BLCW^{+}20, HCR18, MK22]. intelligent [DSL13, TOW^{+}13]. Intent [JA20, LBA^{+}21]. Interacting [LT23, MKH14]. Interaction [Dau18, FSHR^{+}21, Fis22, FN22, FDN20, HCR18, HZ21, HH23, JH18, KKC^{+}22, KLE19, KG18, KF18, LLBN22, LSB^{+}20, ORPW19, Oka18, OH22, PHZ23, PB18, RPB21, RF23, RTC^{+}23, RSB22, SST21, SCK^{+}22, SGHb^{+}23, STH19, SVD21, SZA22, TO21, US23, Vel18, WPC^{+}23, WSG^{+}22, WLTO23, AS17, BFR14, BBR^{+}13, BFL^{+}17, BDO^{+}13, BTLY15, BGJK16, CS13, CXJS13, CBO^{+}16, FSM12, FLA^{+}13, GSF^{+}13, HM13, KSNL13, KTTO13, LMD14, LHP^{+}13, MBK^{+}12, MU15, OP14, RFL^{+}15, dSCV^{+}13, TKC14, THH^{+}13, Wag15, WS13, WBS15, vdBBD^{+}16, DC16, HF14, MF15, SJ18, SJ22]. Interaction-aware [SZA22]. Interactions [HJ23, LO20, NSHK23, ORPW19, RCL^{+}22, SVD21, BWS17, CT12, HMGC13, SFEK15, SDFSM17]. Interactive [AKS21, DRH^{+}23, MLF^{+}21, MD21, SCG17, VPR^{+}22, YZL^{+}16]. interdependence [JBF^{+}14]. interdisciplinary [TGS12]. Interface [CHS^{+}22, LKL23, MBZA20, ZFS^{+}20, PCB^{+}17, RFL^{+}15]. Interfaces [LKL22, DSL13]. Interference [FDN20]. intergroup [DSMS17]. International [GAL^{+}22]. Interpersonal [JDS^{+}21]. interrupt [RFLM13]. Interruptibility [BSC18]. intersection [TGS12]. Intersections [NACB20]. Interventions [SGJ^{+}22]. Intra [BSH^{+}23, GPFA20]. Intra-Hug [BSH^{+}23]. Intra-Robotic-Natural [GPFA20]. Introduction [DC16, DDMZ21, GE13, Goo13, HCR18, HF14, JS19, KB13, MF15, MBZ^{+}22, MAC15, PHZ23, SBB17, TGS12, JS17]. Intuitions [MK22]. Intuitive [LKL22]. invariants [BGJK16]. investigating [WM16]. Investigation [SST21]. invitational [Goo13]. Inviting [Pat23]. ISO [SPGB21]. Issue [DDMZ21, HCR18, JS19, MBZ^{+}22, PHZ23, SBB17, DC16, GE13, Goo13, HF14, KB13, MF15, MAC15, TGS12]. Itself [HPY21]. I’ve [HJL22].

JHRI [Goo13, JS17]. joint [JBF^{+}14]. Journal [SJ22, DC16, HF14, MF15]. Judging [LPPOG23].

Kill [DLG^{+}19]. Kinesthetic

[RSB22, WEG^{+}13]. knowledge [CXJS13].

laboratory [TOW^{+}13]. Lack [vSPKB20]. Language [ECL^{+}22, Ran20, SHSK20]. Laparoscopic [FBGL^{+}22]. law [DC16]. Laying [GPFA20]. Learned [BA18, HJL22, KKC^{+}22]. Learners [ECL^{+}22]. Learning [AOSM^{+}19, AJS19, CNS^{+}20, DLG^{+}19, DKI19, DBK21, FBGL^{+}22, FSLN19, HJL22, KF18, LO20, MLF^{+}21, MNM18, ORPW19, Ran20, SHSK20, SD15, WDR^{+}20, BFL^{+}17, TM12]. led [ECL^{+}22]. legal [JL16]. Legible [BMB^{+}23a, KF21]. Lens [LR23]. Lessons [KKC^{+}22]. Levels [OPK^{+}22, ZZKC19, BFR14]. Light [BMB^{+}23a]. Lightweight [RSB22]. Like [HPY21, SV22]. likeness [MHI16]. Likert [SGHb^{+}23]. Limb [GPFA20]. Linguistic [FSLN19]. Literature [MWJO19, SCK^{+}22, STH19]. Living [Aug14, MNM18]. Load [LLBN22]. Locomotion [LCM^{+}19]. Locus [MvdPD23]. Long [KKC^{+}22, RKD^{+}21, SGJ^{+}22, CBO^{+}16, MN13]. Long-Term [SGJ^{+}22, KKC^{+}22, RKD^{+}21, CBO^{+}16, MN13]. Longer [LRS18]. Longer-Term [LRS18]. looking [WM16]. Loop [BA18, FM22]. Loss [NKT20]. Low [ECL^{+}22].

Low-engaged [ECL^{+}22]. Ludens [MD21].

machines [LZ17, WS13]. magic [Sch14]. Majority [SSQS21]. Make [MK22]. Makers [CK20]. makes [Sch14]. Making [CNS^{+}20, JLV^{+}21, KPSS12, RSB22]. man [WLTO23]. management [SRK14, TKC14]. manipulability [SFEK15]. manipulation [MBK^{+}12]. manipulation-capable [MBK^{+}12]. Manual [GPFA20, dWVK23]. Manufacturer [BMB^{+}23a]. Manufacturing [CHS^{+}22, MBZA20]. Mapping [FGT18]. Matters [FNJB20, RMO23]. may [RFLM13]. Me [SLTK22]. Meal [MN13]. Meal-time [MN13]. Meaning [VWKP22]. measure [LHP^{+}13, MHI16]. Measurement [KNF^{+}22]. Measures [FDN20, MBZA20]. Measuring [BLCW^{+}20, FBGL^{+}22]. mechanical [LJR17, OP14]. Mechanisms [FSLN19]. Media [JS21]. Mediator [NKT20]. Meeting [AKS21]. Melody [FNJB20]. Memory [DKI19, LLBN22]. Mental [HBAZ15, PES20, RKD^{+}21, TdGZ22]. Meta [DWM22]. Meta-analysis [DWM22]. metamorphosis [OP14]. Metaphors [DRH^{+}23]. Method [JDS^{+}21, HLWV14]. Methodology [MWIN22, Tan15]. Methods [BTLY15, LBA^{+}21, MBZ^{+}22, SVD21, TdGZ22, VK19, BDO^{+}13, WS13]. Metrics [MIFP22, MBZA20, NAC^{+}22]. Mind [BKH23, SS18, HJ14]. Minimize [FDC^{+}22]. Minority [SSQS21]. Missing [HY23]. Mixed [BHH^{+}23, CHS21, DLDS^{+}15, SVD21, WPC^{+}23]. Mixed-initiative [CHS21, DLDS^{+}15]. Mobile [BFW^{+}20, CCT^{+}18, CHS21, OPK^{+}22, RKD^{+}21, SPGB21, Vel18, BTLY15, RFL^{+}15]. Mobility [JL19, SD15]. modal [IOLB22]. Modalities [FES22]. Model [CNS^{+}20, DKI19, SST21, WYJ23, CT13, MKH14, MU15, Tan15]. Modeling [SSK^{+}19]. Models [PES20, RKD^{+}21, SZA22, FSM12]. Modified [ZFS^{+}20]. Momentum [MAOTK22]. Monitoring [ME20]. Moral [WKP^{+}23]. Morality [BKH23]. Morally [JW22]. Morphological [KBvdP23]. Morphology [KLE19]. Motion [BA18, FM22, FES22, KF21, LKL23, RPB21, SOCAG^{+}19, MHI16]. Motions [VK19]. Motor [FDN20]. mounted [CHS^{+}22]. movement [HJ14, Tan15]. Movie [Sch14]. Moving [WRPZ21]. mu [MHI16]. Multi [IOLB22, CBO^{+}16, KTTO13, RC15, SFEK15]. multi-activity [CBO^{+}16]. Multi-modal [IOLB22]. multi-party [KTTO13]. multi-robot [SFEK15]. multi-user [RC15]. Multiagent [BK22]. multidisciplinary [BFL^{+}17]. Multilevel [ASJD19]. Multimodal [BBR^{+}13, KLE19, ZCG^{+}20, CT12, LMD14]. Multimodality [LT23]. Multiparty [SSK^{+}19]. Multiplayer [MD21]. Multiple [LT23]. Muscular [FDC^{+}22]. Museum [MLF^{+}21].

Natural [FLA^{+}13, GPFA20]. Navigation [BWS^{+}22, MAOTK22, MBW^{+}23, RVC22]. near [CS13]. near-infrared [CS13]. Need [BMB23b, HPY21, GCSY16]. Needs [BOMLT23, MM16]. Negation [FSLN19]. Negative [CCL^{+}19]. Negotiation [MHV^{+}21, RSB22]. nets [CT12]. Network [BK22, DKI19, GSF^{+}13, PCB^{+}17]. networked [RC15]. Neural [DKI19]. Neural-network-based [DKI19]. Neurophysiological [BK22]. Next [FSHR^{+}21]. No [FSLN19]. Nodding [JAG^{+}23]. Non [SCK^{+}22, DLDS^{+}15, SCG17]. Non-Dyadic [SCK^{+}22]. non-pilots [DLDS^{+}15]. non-technical [SCG17]. Nonanthropomorphic [LCM^{+}19]. Nonverbal [MHV^{+}21, US23]. Norm [SAS19, WKP^{+}23]. Norm-Based [WKP^{+}23]. Normative [SSQS21]. Norms [WKP^{+}23]. Number [WE20]. Nursing [LKL22].

obedience [GCSY16]. Object [FGT18, RTSL23, HS15]. Objective [ME20, RMO23]. objects [LZ17]. Observation [RSB22]. Observed [dWVK23]. off [RPB21]. Older [BFW^{+}20, GAG^{+}21, GAL^{+}22, JL19, PPRR20, MN13]. One [SSQS21]. online [HS15]. Ontology [ORPW19]. Open [IOLB22, SSK^{+}19, SH19, CXJS13]. Open-Source [SH19]. Open-World [SSK^{+}19]. Operated [CHS21]. operation [NWS15]. Operator [ZZKC19, MU15]. Opportunity [Vel18]. Our [AGND23, SJ19b]. Outcomes [RMO23]. Outline [LPPOG23]. overlap [Sch14]. overview [OP14]. Oz [Rie12].

Paediatric [CBK^{+}18]. paired [SD15]. paradigm [KA13]. parametrized [CT13]. Parkinson [BOMLT23, WTDS20]. Participatory [AORK22, GAL^{+}22]. Partner [RSB22, VPR^{+}22]. parts [LJR17]. party [KTTO13]. Passages [BJ23]. Past [Dau18, HY23]. Patient [Fis21]. pedestrian [ITI15]. peer [HBAZ15]. peer-based [HBAZ15]. People [HPY21, JLV^{+}21, SRHE21, WTDS20, BRSTV17, MBK^{+}12, MM16, Wag15]. Perceive [AGND23]. Perceived [BKH23, BLCW^{+}20, CCL^{+}19, WAAR20, MHI16, YZL^{+}16]. Perception [JS21, LPPOG23, LKL23, PT20, SRHE21, TCR20, WE20, vSPKB20]. Perception-Motion [LKL23]. perceptions [WBS15]. Perceptual [TZ21]. Performance [JW21, MBZ^{+}22, HBAZ15, vdBBD^{+}16]. performed [dWVK23]. Performing [AGND23]. Periodically [DBK21]. Personal [LLBN22]. Personality [NKT20, WRPZ21]. Personalization [LRS18]. Personalized [SGJ^{+}22]. Personalizing [MNM18]. Personas [WRPZ21]. Persons [GHD21]. Perspective [BK22, MIFP22, US23, WdHVvR23, KA13, PJ17]. Perspectives [BMB^{+}23a, PHZ23, GE13]. Persuasive [BKH23]. Petri [CT12]. Phase [SST21]. pHRI [FM22]. Physical [AGND23, LO20, ME20, GCMC15, MU15, PCB^{+}17]. Physically [CTA21, CHS^{+}22]. Physiological [ZCG^{+}20]. physiologically [LJR17]. pilots [DLDS^{+}15]. Place [Jør23]. Places [NSHK23]. planar [FSP15]. Planned [PES20]. Planning [BA18, CNS^{+}20, FDC^{+}22, NKFS18]. Plant [BBM^{+}21]. platform [LMD14]. Please [GCSY16]. police [Asa16]. Policies [ORPW19]. policy [DC16]. poor [vdBBD^{+}16]. position [RC15]. Power [BEBT22, TRK21]. Practice [Mat18]. Practices [SGHb^{+}23]. Pre [KF21, RTSL23]. Pre-attentive [KF21]. Pre-beginnings [RTSL23]. Predictability [SRHE21]. Predicting [SVKF23]. Prediction [ZCG^{+}20, MU15]. Preferences [CTA21, KAG21]. Preliminary [JBJ^{+}20, RFLM13]. Presence [MWJO19, NACB20]. Primer [Fis21, HZ21, SVD21]. Principles [KKC^{+}22]. privacy [WM16]. Private [BFW^{+}20]. Probabilistic [JA20]. Problem [PG22, TZ21, BFL^{+}17]. Problem-solving [PG22]. Problematic [HJ23]. Process [JS18]. Processed [WSG^{+}22]. Processing [KF21]. Produce [MK22]. Product [DBK21]. Professional [SV22]. Professionals [BOMLT23]. Proficiencies [LNG22]. Proficiency [NAC^{+}22]. Profit [FNJB20]. progress [MBLD17]. Prohibition [FSLN19]. Project [GAL^{+}22]. projects [GE13]. Promote [LR18, MBLD17, TM12]. Prosthetics [FGCB23]. prototype [SRK14]. proxemic [MM16]. Proximate [MWR22, WAAR20].

Psychological [vSPKB20]. Public [HFK^{+}22, MLF^{+}21, NSHK23, SPGB21, SLTK22, BDO^{+}13, GSF^{+}13]. Publications [SCK^{+}22]. Punish [JW21]. Punishment [JW21]. Purpose [CBK^{+}18, dSCV^{+}13]. push [GCMC15].

Quality [FBGL^{+}22, FN22, FDN20]. Questions [HJL22].

R/E [THH^{+}13]. R2D2 [KLE19]. Rack [RKD^{+}21]. RALL [Ran20]. reactions [LHP^{+}13]. Reactive [BA18]. Real [LO20, BFL^{+}17]. Real-Time [LO20]. real-world [BFL^{+}17]. Reality [BHH^{+}23, CHS^{+}22, HFK^{+}22, WPC^{+}23]. Reasoning [WDR^{+}20]. receive [DSPM15]. receiving [TM12]. Recognition [JA20, NMW^{+}21]. reconnaissance [WML15]. reduce [HLWV14]. redundant [WEG^{+}13]. Reflecting [MWJO19]. reformulation [MBLD17]. Reframing [LR18, MBLD17]. regulating [Har16]. regulation [CT13, LNH23]. Rehabilitation [CBK^{+}18, LLT21, Tan15]. Reinforcement [KF18, MLF^{+}21]. Rejective [FSLN19]. related [Tan15]. Relationship [vSPKB20]. relationships [Mil13]. Release [LSB^{+}20]. Relevance [Mat18]. Reliability [WAAR20]. Reminders [BBM^{+}21]. remote [DSPM15, MBK^{+}12]. Remotely [CHS21]. repertoire [HM13]. Reporting [BEBT22, Rie12]. Representation [AJS19, VWKP22, RFL^{+}15]. Representations [Gär19]. Requests [JW22]. Research [FSHR^{+}21, Jør23, RMO23, KPSS12, YZL^{+}16]. resource [MPCV13]. Results [BFW^{+}20, TOW^{+}13]. resupply [DSPM15]. Reveal [HJL22]. Review [GHD21, JS18, RTC^{+}23, SCK^{+}22, STH19, TdGZ22, WTC^{+}21, dWVK23, AS17, Rie12, ZCV^{+}17]. Reviewer [FS17]. Reviewing [FS17, WLTO23]. Reward [HJL22]. rhythm [MHI16]. rise [LZ17]. roadable [DLDS^{+}15]. Robo [MD21]. Robot [AGND23, AOSM^{+}19, AORK22, BMB^{+}23a, BFW^{+}20, BSC18, BKH23, BBM^{+}21, BSH^{+}23, CBK^{+}18, CHS^{+}22, CCT^{+}18, CNS^{+}20, CK20, DC16, Dau18, DRH^{+}23, DLG^{+}19, DKI19, DBK21, ESK23, ECL^{+}22, FSHR^{+}21, FBGL^{+}22, FDC^{+}22, Fis22, FN22, FES22, Gär19, GAG^{+}21, HGA^{+}21, HPY21, HY23, HCR18, HHHA19, HFK^{+}22, HZ21, HF14, HH23, JAG^{+}23, JW21, JH18, JDS^{+}21, KKC^{+}22, KLE19, KG18, KF21, KBvdP23, LLBN22, LBA^{+}21, LRS18, LCM^{+}19, LPPOG23, LNH23, LKL22, LNG22, LO20, LSB^{+}20, MIFP22, MF15, MBZ^{+}22, MWR22, MAOTK22, MBW^{+}23, MTCKS21, ME20, MvdPD23, MHV^{+}21, MNM18, NKFS18, NKT20, NAC^{+}22, Oka18, OPK^{+}22, OH22, PG22, Pat23, PHZ23, PT20, PPRR20, PB18, RPB21, RF23, RTC^{+}23, RVC22, RTSL23, SJ18, SJ22, SST21, SRHE21, SCK^{+}22, STH19, SVKF23, SGJ^{+}22, SH21, SH19, Tan15]. Robot [TCR20, TO21, US23, Vel18, VK19, VPR^{+}22, WSG^{+}22, WAAR20, WKP^{+}23, WDR^{+}20, WTDS20, WLTO23, WE20, ZCG^{+}20, dWVK23, vSPKB20, AS17, BRSTV17, BFR14, BBR^{+}13, BFL^{+}17, BDO^{+}13, BTLY15, CS13, CXJS13, CBO^{+}16, DTT^{+}13, DSMS17, FM13, FSM12, FLA^{+}13, GSF^{+}13, GCMC15, HBAZ15, HMGC13, HS15, HM13, KA13, KTTO13, LMD14, LHP^{+}13, LJR17, MHI16, MN13, MM16, MPCV13, MKH14, MU15, PSB^{+}16, PJ17, RFL^{+}15, dSCV^{+}13, SRK14, Sch14, SGHb^{+}23, SVD21, SFEK15, SA15, SLD^{+}13, TM12, TKC14, THH^{+}13, Wag15, WPC^{+}23, WBS15, YZL^{+}16, vdBBD^{+}16, Ran20]. Robot-AI [Vel18]. Robot-aided [Tan15]. Robot-Assisted [JDS^{+}21, ZCG^{+}20, Ran20]. Robot-Centric [TCR20]. Robot-Delivered [GAG^{+}21]. Robot-Initiated [JW21]. Robot-Interaction [PB18]. Robot-led [ECL^{+}22]. Robot-performed [dWVK23]. Robot-specific [KLE19]. Robot-to-Human [FES22, PJ17]. Robotic [AKS21, DDMZ21, GPFA20, JW22, JBJ^{+}20, JS21, KDC^{+}21, MWIN22, RTC^{+}23, SOCAG^{+}19, OP14]. Robotics [CTA21, CK20, FGCB23, HB18, JA20, Jør23, TZ21, YZL^{+}16, ZCV^{+}17]. Robots [AGND23, BOMLT23, BLCW^{+}20, BJ23, BSH^{+}23, BMB23b, CCL^{+}19, CHS21, FNJB20, FSLN19, GAL^{+}22, GHD21, GRSA23, JW21, JLV^{+}21, Jør23, JH18, KAG21, KBvdP23, LLT21, LCTRS22, LR18, LR23, MM16, MK22, MWJO19, MD21, NSHK23, SSQS21, SPGB21, SS18, SLTK22, SHSK20, SOG^{+}22, TdGZ22, Vel18, Wag15, WRPZ21, Aug14, BWS17, BTLY15, GCSY16, Har16, HJ14, HM13, HLWV14, MBK^{+}12, SRK14, SCG17, Sch14, SDFSM17, TOW^{+}13, WML15, WEG^{+}13]. Robots-from [SS18]. Robust [JH18]. Role [SAS19, WKP^{+}23, HLWV14]. Role-Based [WKP^{+}23]. Role-Grounded [WKP^{+}23]. Roles [LT23, LLT21, SV22, HLWV14]. Room [LLBN22]. Roomba [JL16]. Route [RVC22]. Routes [PES20].

Safety [SPGB21]. Satisfaction [JAG^{+}23]. Say [FSLN19]. Scaffold [SHSK20]. Scaffolding [PG22]. Scalable [FDN20]. Scale [PB18, SGHb^{+}23]. Scales [ASJD19]. Scenario [FN22]. Scenarios [NMW^{+}21]. Scholarship [SJ19b]. Science [Jør23, KG18, MWJO19]. sciences [TGS12]. scientific [MBLD17]. Script [BKH23]. SE [PB18]. SE-HRI [PB18]. seamless [SLD^{+}13]. Second [ECL^{+}22, SHSK20]. Second-language [SHSK20]. See [SRHE21]. Seeking [BKH23]. Selection [CRA21]. Self [NKT20, NAC^{+}22, PB18, WYJ23]. Self-assessment [NAC^{+}22]. Self-confidence [WYJ23]. Self-Disclosure [NKT20]. Self-Efficacy [PB18]. Semantics [Gär19]. Sensitive [JW22]. Sensory [JBJ^{+}20]. Service [BFW^{+}20, Vel18]. Serving [SPGB21]. Shaping [SJ19b]. Shared [CHS^{+}22, FM22, FN22, JA20, LBA^{+}21, MAC15, NWS15, SA15, SD15]. Shoe [RKD^{+}21]. shoot [Asa16]. Shopkeeper [DLG^{+}19, DBK21]. Shopworker [ESK23]. Should [JLV^{+}21, BRSTV17]. side [MKH14]. side-by-side [MKH14]. signal [MM16]. Signals [ZCG^{+}20, vdBBD^{+}16]. Simulation [BWS^{+}22]. Situ [CBK^{+}18, SRK14]. situation [PJ17]. Size [LLBN22]. Skills [PG22]. Skin [HH23]. smart [SD15]. smoothness [Tan15]. Social [AS17, AORK22, BKH23, BLCW^{+}20, BWS^{+}22, CBK^{+}18, DKI19, GHD21, HH23, KAG21, LLT21, Lju23, MAOTK22, MBW^{+}23, NKT20, PT20, RMO23, SRHE21, SLTK22, SHSK20, SVD21, SV22, SH21, TZ21, TO21, WTC^{+}21, WRPZ21, WDR^{+}20, BBR^{+}13, CT13, CBO^{+}16, GSF^{+}13, Goo13, HM13, MM16, dSCV^{+}13, Sch14, You17]. Socially [BOMLT23, DRH^{+}23, LPPOG23, MAOTK22, MNM18, MD21, SGJ^{+}22, SOG^{+}22, WTDS20, FM13, HLWV14, MN13, SDFSM17, TOW^{+}13]. SocNavBench [BWS^{+}22]. Soft [Jør23]. solving [BFL^{+}17, PG22]. Some [Dau18]. Sorry [RFLM13]. sound [RFL^{+}15]. Source [SH19, RFL^{+}15]. Space [LLBN22, WEG^{+}13]. Spaces [SPGB21, SLTK22, GSF^{+}13]. Spatial [PES20, RPB21]. Speaking [WRPZ21]. Special [DDMZ21, HCR18, MBZ^{+}22, PHZ23, SBB17, DC16, GE13, Goo13, HF14, KB13, MF15, MAC15, Sch14, TGS12]. Specific [BOMLT23, KLE19]. spectroscopy [CS13]. Spectrum [JBJ^{+}20]. speculative [Aug14]. Speech [FNJB20, NMW^{+}21, BWS17]. Spring [SJ19b]. standard [SPGB21]. Standards [AKS21]. State [SJ22, TdGZ22]. Statements [JAG^{+}23]. States [PT20].

Statistical [BEBT22]. status [NWS15]. steering [NWS15]. stereotype [Wag15]. stereotypes [HLWV14]. stiffness [MU15]. Stimulation [GAG^{+}21]. Stochastic [LNG22]. Stop [SLTK22]. Storytellers [SOG^{+}22]. Strategies [RPB21, WKP^{+}23, ZZKC19]. Structure [LPPOG23]. students [ZCV^{+}17]. Studies [KDC^{+}21, STH19, US23, KB13, LMD14, Rie12, TKC14, TOW^{+}13]. Study [AOSM^{+}19, BSC18, CTA21, FBGL^{+}22, FES22, HFK^{+}22, JBJ^{+}20, JAG^{+}23, JDS^{+}21, KAG21, MLF^{+}21, RSB22, RKD^{+}21, SHSK20, KSNL13, WEG^{+}13, ZCW^{+}16]. Subjective [RMO23]. subjects [SCG17]. Subterranean [RF23]. Successful [LR18]. Supervisory [HHHA19, Mil13]. support [JBF^{+}14, YZL^{+}16]. Supporting [PJ17]. suppression [MHI16]. Surgery [FBGL^{+}22, ZCG^{+}20, ZCW^{+}16]. surroundings [PCB^{+}17]. Survey [MBW^{+}23, Ran20, RTC^{+}23, VK19, WPC^{+}23]. sustainability [JS17]. Swarm [KF21, WdHVvR23, BGJK16, KSNL13]. Swarms [SOCAG^{+}19, KSNL13, SFEK15]. switching [CBO^{+}16]. System [DLG^{+}19, KKC^{+}22, BDO^{+}13, GSF^{+}13, ITI15, KB13, KTTO13, NWS15, TKC14]. Systematic [GHD21, RTC^{+}23, STH19, TdGZ22, dWVK23, Rie12]. Systems [AKS21, DDMZ21, MLF^{+}21, SA15, TGS12].

Tactile [BJ23, JS21, LJR17]. Tactile-Affects [BJ23]. Takes [LNH23]. taking [CT12, PJ17]. Tale [BAC^{+}23]. Talking [WDR^{+}20]. Tangled [BJ23]. Tap [GCMC15]. Target [CRA21]. Task [FGT18, JW21, LNG22, VWKP22, VPR^{+}22, BWS17, HBAZ15, Tan15, WEG^{+}13]. task-based [BWS17]. task-related [Tan15]. Tasks [AGND23, BK22, BA18, CHS^{+}22, MWIN22, RSB22, GCMC15]. Taxonomy [MD21, TO21, WPC^{+}23]. teach [TM12]. Teacher [AGND23]. Teaching [VPR^{+}22, SCG17, WEG^{+}13]. Team [MIFP22, RF23]. Teaming [MWIN22, MBZ^{+}22]. Teams [CHS^{+}22, HHHA19, KNF^{+}22, LNG22, MIFP22, NAC^{+}22, HBAZ15, PSB^{+}16]. Teamwork [MIFP22]. technical [Goo13, SCG17]. Technique [ZFS^{+}20]. Techniques [STH19]. technological [HLWV14]. technologists [You17]. Technology [JL19, JL16, You17]. Tele [LKL22]. Tele-Nursing [LKL22]. Teleoperated [JS21, OPK^{+}22]. Teleoperation [CCT^{+}18, DSL13, LKL23, NACB20, PCB^{+}17, RFL^{+}15, WS13]. Telepresence [LKL23, HMGC13]. telerobotic [ZCW^{+}16]. Telerobots [BAC^{+}23]. Term [LRS18, SGJ^{+}22, CBO^{+}16, KKC^{+}22, MN13, RKD^{+}21]. Test [MBZ^{+}22]. Testing [BWS^{+}22, GAL^{+}22, LMD14]. Texture [HH23]. Texture-changing [HH23]. Their [NACB20, MM16, TM12]. theoretic [DTT^{+}13]. Theories [JH18]. Theory [Mat18, TRK21]. Therapy [BBM^{+}21, BMB23b]. There [SRHE21]. Thermal [PT20]. These [WM16]. Think [JL19, PES20]. thinking [MBLD17]. thinks [JL16]. Thoughts [Dau18]. threat [YZL^{+}16]. Three [FS17]. THRI [JS18, JS19]. Time [AKS21, ASJD19, JH18, LO20, MN13]. timed [CT12]. Timing [CT12]. Together [FGCB23, MKH14]. too [MM16]. Toolkit [KNF^{+}22]. Touch [Pat23]. Touching [LJR17]. Tough [TZ21]. Tower [JDS^{+}21]. Tracking [Fis22]. Traditional [JL19]. Trainee [AGND23]. Training [ZZKC19]. Traits [NKT20]. Trajectory [KF21, LO20, Tan15]. Transactions [SJ18, SJ22]. Transfer [FGT18]. Transparency [vSPKB20, RCL^{+}22]. Transport [HFK^{+}22]. Trees [HGA^{+}21]. Trends [SGHb^{+}23]. Trial [ESK23]. Trials [BFW^{+}20, TOW^{+}13]. Trivialization [SLTK22]. Trust [AGND23, BKH23, CNS^{+}20, JAG^{+}23, KNF^{+}22, OH22, WAAR20, WYJ23, DSMS17, PSB^{+}16]. Trust-Aware [CNS^{+}20]. trustworthy [ITI15]. tu [Har16]. turn [CT12]. turn-taking [CT12]. Tutoring [LRS18]. Tutors [SGJ^{+}22]. Two [BAC^{+}23, BGJK16, LNH23, BRSTV17, KSNL13, ZCV^{+}17]. types [KSNL13].

UAV [ZFS^{+}20]. UGV [PCB^{+}17]. UMRF [VWKP22]. Uncanny [DWM22]. Unconventional [ZCV^{+}17]. Understand [Lju23, NSHK23, BWS17]. Understanding [DRH^{+}23, JS18, KAG21, SRHE21, SDFSM17, MM16, MTCKS21]. Unguided [NSHK23]. Unified [VWKP22]. uninhabited [Mil13]. unmanned [DSPM15]. Usability [GAG^{+}21, KDC^{+}21]. Usage [SGHb^{+}23]. Use [CCL^{+}19, Asa16, BTLY15]. Useful [CTA21, KPSS12]. User [CTA21, DRH^{+}23, IOLB22, LPPOG23, MBK^{+}12, NSHK23, STH19, SVKF23, WML15, RC15, RFL^{+}15, WEG^{+}13]. User-centered [MBK^{+}12, WML15]. User-Generated [NSHK23]. Users [BAC^{+}23, CBO^{+}16]. Using [FNJB20, Gär19, HGA^{+}21, LCM^{+}19, LNH23, ME20, MD21, NSHK23, CT12, CBO^{+}16, SFEK15, Tan15, Wag15, ZCW^{+}16].

Validating [SZA22]. Validation [PB18]. Valley [DWM22]. value [GCMC15]. Variable [CHS21]. Variables [DWM22]. vehicle [DLDS^{+}15]. Vehicles [SZA22, Mil13]. Verbal [HPY21, NKFS18]. Via [FES22, CCT^{+}18, FN22, PT20]. Viable [GPFA20]. Video [JAG^{+}23]. Videos [NSHK23, WM16]. Viewed [MK22]. Virtual [HFK^{+}22, RSB22, WPC^{+}23, WS13]. Vision [KDC^{+}21, RTC^{+}23]. Vision-Based [KDC^{+}21]. Visual [RFL^{+}15, SSK^{+}19]. visuospatial [PJ17]. vocabulary [TM12].

Volume [JS19].

Walking [MKH14]. Warning [vdBBD^{+}16]. Welcome [SJ18]. We’re [WDR^{+}20]. Wheelchair [KDC^{+}21]. Where [FSHR^{+}21, PES20]. whether [MK22]. Who [WLTO23]. Wild [JH18, MvdPD23, RKD^{+}21, SRK14]. Wisdom [LR23]. within [JS21]. Wizard [Rie12]. Work [LLBN22, TGS12]. Worker [MTCKS21]. Working [LLBN22]. Workload [HHHA19, ME20, ZCG^{+}20, HBAZ15]. Workspaces [LBA^{+}21]. World [IOLB22, SSK^{+}19, BFL^{+}17]. Worlds [BHH^{+}23]. Would [HPY21].

X [SVD21].

year [GAL^{+}22]. Years [SCK^{+}22, WLTO23]. Young [BBM^{+}21, CBO^{+}16]. You’re [PES20, WM16]. YouTube [NSHK23].

Zhao [Fis21, SVD21].

References

Aliasghari:2023:HDW

[AGND23] Pourya Aliasghari, Moojan Ghafurian, Chrystopher L. Nehaniv, and Kerstin Dautenhahn. How do we perceive our trainee robots? Exploring the impact of robot errors and appearance when performing domestic physical tasks on teachers’ trust and evaluations. ACM Transactions on Human– Robot Interaction (THRI), 12 (3):37:1–37:??, September 2023. CODEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3582516.

Aly:2019:ERL

[AJS19] Amir Aly, Odest Chadwicke Jenkins, and Selma Sabanovic. Editorial: Representation learning in HRI. ACM Transactions on Human–Robot Interaction (THRI), 8(4):19:1–19:2, December 2019. CODEN ???? ISSN 2573-9522 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3366621.

Arnold:2021:ETM

[AKS21] Thomas Arnold, Daniel Kasenberg, and Matthias Scheutz. Explaining in time: Meeting interactive standards of explanation for robotic systems. ACM Transactions on Human– Robot Interaction (THRI), 10 (3):25:1–25:23, July 2021. CO-DEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3457183.

Axelsson:2022:SRC

[AORK22] Minja Axelsson, Raquel Oliveira, Mattia Racca, and Ville Kyrki. Social robot co-design canvases: a participatory design framework. ACM Transactions on Human–Robot Interaction (THRI), 11(1):3:1–3:39, March 2022. CODEN ???? ISSN 2573-9522 (electronic). URL

https://dl.acm.org/doi/10.
1145/3472225.

Alves-Oliveira:2019:ERG

[AOSM^{+}19] Patrícia Alves-Oliveira, Pedro Sequeira, Francisco S. Melo, Ginevra Castellano, and Ana Paiva. Empathic robot for group learning: a field study. ACM Transactions on Human– Robot Interaction (THRI), 8(1): 3:1–3:34, March 2019. CO-DEN ???? ISSN 2573-9522 (electronic). URL https://dl.

                                        acm.org/doi/abs/10.1145/
                                        3300188.

Admoni:2017:SEG

[AS17] Henny Admoni and Brian Scassellati. Social eye gaze in human–robot interaction: a review. Journal of Human–Robot Interaction (JHRI), 6(1):25– 63, May 2017. ISSN 2163-

                                        0364. URL https://dl.acm.
                                        org/doi/abs/10.5898/JHRI.
                                        6.1.Admoni.

Asaro:2016:HDS

[Asa16] Peter Asaro. “Hands up, don’t shoot!”: HRI and the automation of police use of force. Journal of Human–Robot Interaction (JHRI), 5(3):55–69, December 2016. ISSN 2163-

                                        0364. URL https://dl.acm.
                                        org/doi/abs/10.5898/JHRI.
                                        5.3.Asaro.

Asselborn:2019:BMT

[ASJD19] Thibault Asselborn, Kshitij Sharma, Wafa Johal, and Pierre Dillenbourg. Bridging multilevel time scales in HRI: an analysis framework. ACM Transactions on Human–Robot Interaction (THRI), 8(3):17:1– 17:24, August 2019. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3338809.

Auger:2014:LRS

[Aug14] James Auger. Living with robots: a speculative design approach. Journal of Human– Robot Interaction (JHRI), 3(1): 20–42, February 2014. ISSN 2163-0364. URL https://dl.

acm.org/doi/abs/10.5898/
JHRI.3.1.Auger.

Bowen:2018:CLG

[BA18] Chris Bowen and Ron Alterovitz. Closed-loop global motion planning for reactive, collision-free execution of learned tasks. ACM Transactions on Human–Robot Interaction (THRI), 7(1):10:1– 10:16, May 2018. CODEN ???? ISSN 2573-9522 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3209045.

Benson:2023:EEG

[BAC^{+}23] William N. Benson, Zachary Anderson, Evan Capelle, Maya F. Dunlap, Blake Dorris, Jenna L. Gorlewicz, Mitsuru Shimizu, and Jerry B. Weinberg. Embodied expressive gestures in telerobots: a tale of two users. ACM Transactions on Human– Robot Interaction (THRI), 12 (2):24:1–24:??, June 2023. CO-DEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3570908.

Bhat:2021:PRH

[BBM^{+}21] Ashwin Sadananda Bhat, Christiaan Boersma, Max Jan Meijer, Maaike Dokter, Ernst Bohlmeijer, and Jamy Li. Plant robot for at-home behavioral activation therapy reminders to young adults with depression. ACM Transactions on Human– Robot Interaction (THRI), 10 (3):20:1–20:21, July 2021. CO-DEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3442680.

Belpaeme:2013:MCR

[BBR^{+}13] Tony Belpaeme, Paul Baxter, Robin Read, Rachel Wood, Heriberto Cuayáhuitl, Bernd Kiefer, Stefania Racioppa, Ivana Kruijff-Korbayová, Georgios Athanasopoulos, Valentin Enescu, Rosemarijn Looije, Mark Neerincx, Yiannis Demiris, Raquel Ros-Espinoza, Aryel Beck, Lola Cañamero, Antione Hiolle, Matthew Lewis, Ilaria Baroni, Marco Nalin, Piero Cosi, Giulio Paci, Fabio Tesser, Giacomo Sommavilla, and Remi Humbert. Multimodal child–robot interaction: building social bonds. Journal of Human–Robot Interaction (JHRI), 1(2):33–53, February 2013. ISSN 2163-0364.

URL https://dl.acm.org/
doi/abs/10.5555/3109688.
3109691.

Breazeal:2013:CHR

[BDO^{+}13] Cynthia Breazeal, Nick De-Palma, Jeff Orkin, Sonia Chernova, and Malte Jung. Crowdsourcing human–robot interaction: new methods and system evaluation in a public environment. Journal of Human–Robot Interaction (JHRI), 2(1):82– 111, February 2013. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
2.1.Breazeal.

Bartlett:2022:GPA

[BEBT22] Madeleine E. Bartlett, C. E. R.

Edmunds, Tony Belpaeme, and Serge Thill. Have I got the power? Analysing and reporting statistical power in HRI. ACM Transactions on Human– Robot Interaction (THRI), 11 (2):16:1–16:??, June 2022. CO-DEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3495246.

Blain:2017:MAL

[BFL^{+}17] Rob Blain, Alex Ferworn, Jean Li, Jimmy Tran, and Michael Carter. A multidisciplinary approach to learning human–robot interaction (HRI) through real-world problem solving: the “BUSA Dig”. Journal of Human–Robot Interaction (JHRI), 6(2):70–91, September 2017. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
6.2.Blain.

Beer:2014:TFL

[BFR14] Jenay M. Beer, Arthur D. Fisk, and Wendy A. Rogers. Toward a framework for levels of robot autonomy in human–robot interaction. Journal of Human– Robot Interaction (JHRI), 3(2): 74–99, July 2014. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
3.2.Beer.

Bajones:2020:RFT

[BFW^{+}20] Markus Bajones, David Fischinger, Astrid Weiss, Paloma De La Puente, Daniel Wolf, Markus Vincze, Tobias Körtner, Markus Weninger, Konstantinos Papoutsakis, Damien Michel, Ammar Qammaz, Paschalis Panteleris, Michalis Foukarakis, Ilia Adami, Danae Ioannidi, Asterios Leonidis, Margherita Antona, Antonis Argyros, Peter Mayer, Paul Panek, Håkan Eftring, and Susanne Frennert. Results of field trials with a mobile service robot for older adults in 16 private households. ACM Transactions on Human–Robot Interaction (THRI), 9(2):10:1–10:27, February 2020. CODEN ???? ISSN 2573-9522 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3368554.

Brown:2016:TIH

[BGJK16] Daniel S. Brown, Michael A. Goodrich, Shin-Young Jung, and Sean Kerman. Two invariants of human–swarm interaction. Journal of Human– Robot Interaction (JHRI), 5(1): 1–31, March 2016. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
5.1.Brown.

Brown:2023:BBW

[BHH^{+}23] Landon Brown, Jared Hamilton, Zhao Han, Albert Phan, Thao Phung, Eric Hansen, Nhan Tran, and Tom Williams. Best of both worlds? Combining different forms of mixed reality deictic gestures. ACM Transactions on Human–Robot Interaction (THRI), 12(1):9:1– 9:??, March 2023. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3563387.

Barker:2023:CRT

[BJ23] Ned Barker and Carey Jewitt.

Collaborative robots and tangled passages of tactile-affects. ACM Transactions on Human– Robot Interaction (THRI), 12 (2):19:1–19:??, June 2023. CO-DEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3534090.

Bales:2022:NBD

[BK22] Gregory Bales and Zhaodan Kong. Neurophysiological and behavioral differences in human-multiagent tasks: an EEG network perspective. ACM Transactions on Human– Robot Interaction (THRI), 11 (4):42:1–42:??, December 2022. CODEN ???? ISSN 2573-9522 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3527928.

Banks:2023:AAS

[BKH23] Jaime Banks, Kevin Koban, and Brad Haggadone. Avoiding the abject and seeking the script: Perceived mind, morality, and trust in a persuasive social robot. ACM Transactions on Human–Robot Interaction (THRI), 12(3):32:1–32:??, September 2023. CODEN ???? ISSN 2573-9522 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3572036.

Barchard:2020:MPS

[BLCW^{+}20] Kimberly A. Barchard, Leiszle Lapping-Carr, R. Shane Westfall, Andrea Fink-Armold, Santosh Balajee Banisetty, and David Feil-Seifer. Measuring the perceived social intelligence of robots. ACM Transactions on Human–Robot Interaction (THRI), 9(4):24:1– 24:29, October 2020. CODEN ???? ISSN 2573-9522 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3415139.

Bacula:2023:IRM

[BMB^{+}23a] Alexandra Bacula, Jason Mercer, Jaden Berger, Julie Adams, and Heather Knight. Integrating robot manufacturer perspectives into legible factory robot light communications. ACM Transactions on Human– Robot Interaction (THRI), 12 (1):13:1–13:??, March 2023. CODEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3570732.

Bucci:2023:ARN

[BMB23b] Paul Bucci, David Marino, and Ivan Beschastnikh. Affective robots need therapy. ACM Transactions on Human– Robot Interaction (THRI), 12 (2):20:1–20:??, June 2023. CO-DEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3543514.

Bar-On:2023:SAR

[BOMLT23] Inbal Bar-On, Gili Mayo, and Shelly Levy-Tzedek. Socially assistive robots for Parkinson’s disease: Needs, attitudes and specific applications as identified by healthcare professionals. ACM Transactions on Human– Robot Interaction (THRI), 12 (1):11:1–11:??, March 2023. CODEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3570168.

Brooks:2018:BCH

[Bro18] Rodney Brooks. A brave, creative, and happy HRI. ACM Transactions on Human–Robot Interaction (THRI), 7(1):1:1– 1:3, May 2018. CODEN ???? ISSN 2573-9522 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3209540.

Ball:2017:HSR

[BRSTV17] Adrian Keith Ball, David C.

Rye, David Silvera-Tawil, and Mari Velonaki. How should a robot approach two people? Journal of Human–Robot Interaction (JHRI), 6(3):71–91, December 2017. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
6.3.Ball.

Banerjee:2018:RCH

[BSC18] Siddhartha Banerjee, Andrew Silva, and Sonia Chernova. Robot classification of human interruptibility and a study of its effects. ACM Transactions on Human–Robot Interaction (THRI), 7(2):14:1–14:35, October 2018. CODEN ???? ISSN 2573-9522 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3277902.

Block:2023:ARD

[BSH^{+}23] Alexis E. Block, Hasti Seifi, Otmar Hilliges, Roger Gassert, and Katherine J. Kuchenbecker. In the arms of a robot: Designing autonomous hugging robots with intra-hug gestures. ACM Transactions on Human–Robot Interaction (THRI), 12(2):18:1–18:??, June 2023. CODEN ???? ISSN 2573-9522 (electronic). URL

https://dl.acm.org/doi/10.
1145/3526110.

Brooks:2015:MEC

[BTLY15] Daniel J. Brooks, Katherine M. Tsui, Michael Lunderville, and Holly A. Yanco. Methods for evaluating and comparing the use of haptic feedback in human–robot interaction with ground-based mobile robots. Journal of Human– Robot Interaction (JHRI), 4(1): 3–29, July 2015. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
4.1.Brooks.

Briggs:2017:ERU

[BWS17] Gordon Briggs, Tom Williams, and Matthias Scheutz. Enabling robots to understand indirect speech acts in taskbased interactions. Journal of Human–Robot Interaction (JHRI), 6(1):64–94, May 2017. ISSN 2163-0364.

URL https://dl.acm.org/
doi/abs/10.5898/JHRI.6.1.
Briggs.

Biswas:2022:SGS

[BWS^{+}22] Abhijat Biswas, Allan Wang, Gustavo Silvera, Aaron Steinfeld, and Henny Admoni. Soc-NavBench: a grounded simulation testing framework for evaluating social navigation. ACM Transactions on Human– Robot Interaction (THRI), 11 (3):26:1–26:??, September 2022. CODEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3476413.

Carrillo:2018:AGP

[CBK^{+}18] Felip Martí Carrillo, Joanna Butchart, Sarah Knight, Adam Scheinberg, Lisa Wise, Leon Sterling, and Chris McCarthy. Adapting a general-purpose social robot for paediatric rehabilitation through in situ design.

ACM Transactions on Human– Robot Interaction (THRI), 7(1): 12:1–12:30, May 2018. CO-DEN ???? ISSN 2573-9522 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/
3203304.

Coninx:2016:TLT

[CBO^{+}16] Alexandre Coninx, Paul Baxter, Elettra Oleari, Sara Bellini, Bert Bierman, Olivier Blanson Henkemans, Lola Cañamero, Piero Cosi, Valentin Enescu, Raquel Ros Espinoza, Antoine Hiolle, Rémi Humbert, Bernd Kiefer, Ivana Kruijff-Korbayová, Rose marijn Looije, Marco Mosconi, Mark Neerincx, Giulio Paci, Georgios Patsis, Clara Pozzi, Francesca Sacchitelli, Hichem Sahli, Alberto Sanna, Giacomo Sommavilla, Fabio Tesser, Yiannis Demiris, and Tony Belpaeme. Towards long-term social child–robot interaction: using multi-activity switching to engage young users. Journal of Human–Robot Interaction (JHRI), 5(1):32– 67, March 2016. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
5.1.Coninx.

Chien:2019:ADP

[CCL^{+}19] Sung-En Chien, Li Chu, Hsing-Hao Lee, Chien-Chun Yang, Fo-Hui Lin, Pei-Ling Yang, Te-Mei Wang, and Su-Ling Yeh. Age difference in perceived ease of use, curiosity, and implicit negative attitude toward robots. ACM Transactions on Human–Robot Interaction (THRI), 8(2):9:1–9:19, June 2019. CODEN ???? ISSN 2573-9522 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3311788.

Che:2018:FHM

[CCT^{+}18] Yuhang Che, Heather Culbertson, Chih-Wei Tang, Sudipto Aich, and Allison M. Okamura. Facilitating humanmobile robot communication via haptic feedback and gesture teleoperation. ACM Transactions on Human–Robot Interaction (THRI), 7(3):20:1–20:23, December 2018. CODEN ???? ISSN 2573-9522 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3243503.

Chiou:2021:MIV

[CHS21] Manolis Chiou, Nick Hawes, and Rustam Stolkin. Mixedinitiative variable autonomy for remotely operated mobile robots. ACM Transactions on Human–Robot Interaction (THRI), 10(4):37:1–37:34, December 2021. CODEN ???? ISSN 2573-9522 (electronic).

URL https://dl.acm.org/
doi/10.1145/3472206.

Chan:2022:DEA

[CHS^{+}22] Wesley P. Chan, Geoffrey Hanks, Maram Sakr, Haomiao Zhang, Tiger Zuo, H. F. Machiel van der Loos, and Elizabeth Croft. Design and evaluation of an augmented reality headmounted display interface for human robot teams collaborating in physically shared manufacturing tasks. ACM Transactions on Human–Robot Interaction (THRI), 11(3):31:1–31:??, September 2022. CODEN ???? ISSN 2573-9522 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3524082.

Chun:2020:RME

[CK20] Bohkyung Chun and Heather Knight. The robot makers: an ethnography of anthropomorphism at a robotics company. ACM Transactions on Human– Robot Interaction (THRI), 9(3): 16:1–16:36, July 2020. CODEN ???? ISSN 2573-9522 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3377343.

Chen:2020:TAD

[CNS^{+}20] Min Chen, Stefanos Nikolaidis, Harold Soh, David Hsu, and Siddhartha Srinivasa. Trust-aware decision making for human–robot collaboration: Model learning and planning. ACM Transactions on Human–Robot Interaction (THRI), 9(2):9:1–9:23, February 2020. CODEN ???? ISSN 2573-9522 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3359616.

Cody:2021:HCC

[CRA21] Jason R. Cody, Karina A. Roundtree, and Julie A. Adams. Human-collective collaborative target selection. ACM Transactions on Human– Robot Interaction (THRI), 10 (2):18:1–18:29, May 2021. CO-DEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3442679.

Canning:2013:FNI

[CS13] Cody Canning and Matthias Scheutz. Functional nearinfrared spectroscopy in human– robot interaction. Journal of Human–Robot Interaction (JHRI), 2(3):62–84, September 2013. ISSN 2163-0364.

URL https://dl.acm.org/
doi/abs/10.5898/JHRI.2.3.
Canning.

Chao:2012:TMT

[CT12] Crystal Chao and Andrea L.

Thomaz. Timing in multimodal turn-taking interactions: control and analysis using timed Petri nets. Journal of Human– Robot Interaction (JHRI), 1(1): 4–25, July 2012. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
1.1.Chao.

Chao:2013:CSD

[CT13] Crystal Chao and Andrea L.

Thomaz. Controlling social dynamics with a parametrized model of floor regulation. Journal of Human–Robot Interaction (JHRI), 2(1):4–29, February 2013. ISSN 2163-0364.

URL https://dl.acm.org/
doi/abs/10.5898/JHRI.2.1.
Chao.

Canal:2021:PUB

[CTA21] Gerard Canal, Carme Torras, and Guillem Alenyà. Are preferences useful for better assistance?: a physically assistive robotics user study. ACM Transactions on Human– Robot Interaction (THRI), 10 (4):39:1–39:19, December 2021. CODEN ???? ISSN 2573-9522 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3472208.

Chen:2013:TOK

[CXJS13] Xiaoping Chen, Jiongkun Xie, Jianmin Ji, and Zhiqiang Sui. Toward open knowledge enabling for human–robot interaction. Journal of Human–Robot Interaction (JHRI), 1(2):100– 117, February 2013. ISSN 2163-

                                        0364. URL https://dl.acm.
                                        org/doi/abs/10.5898/JHRI.
                                        1.2.Chen.

Dautenhahn:2018:SBT

[Dau18] Kerstin Dautenhahn. Some brief thoughts on the past and future of human–robot interaction. ACM Transactions on Human–Robot Interaction (THRI), 7(1):4:1–4:3, May 2018. CODEN ???? ISSN 2573-9522 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3209769.

Doering:2021:DDI

[DBK21] Malcolm Doering, Drazen Brscić, and Takayuki Kanda. Data-driven imitation learning for a shopkeeper robot with periodically changing product information. ACM Transactions on Human–Robot Interaction (THRI), 10(4):31:1– 31:20, December 2021. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3451883.

Darling:2016:IJH

[DC16] Kate Darling and Ryan Calo.

Introduction to Journal of Human–Robot Interaction special issue on law and policy. Journal of Human–Robot Interaction (JHRI), 5(3):1–2, December 2016. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
5.3.Darling.

DeGraaf:2021:ISI

[DDMZ21] Maartje M. A. De Graaf, Anca Dragan, Bertram F. Malle, and Tom Ziemke. Introduction to the special issue on explainable robotic systems. ACM Transactions on Human–Robot Interaction (THRI), 10(3):22:1– 22:4, July 2021. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3461597.

Doering:2019:NNB

[DKI19] Malcolm Doering, Takayuki Kanda, and Hiroshi Ishiguro. Neural-network-based memory for a social robot: Learning a memory model of human behavior from data. ACM Transactions on Human–Robot Interaction (THRI), 8(4):24:1– 24:27, December 2019. CO-DEN ???? ISSN 2573-9522 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/
3338810.

Dorneich:2015:MIC

[DLDS^{+}15] Michael C. Dorneich, Emmanuel Letsu-Dake, Sanjiv Singh, Sebastian Scherer, Lyle Chamberlain, and Marcel Bergerman. Mixed-initiative control of a roadable air vehicle for nonpilots. Journal of Human–Robot Interaction (JHRI), 4(3):38–61, December 2015. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
4.3.Dorneich.

Doering:2019:CDK

[DLG^{+}19] Malcolm Doering, Phoebe Liu, Dylan F. Glas, Takayuki Kanda, Dana Kulić, and Hiroshi Ishiguro. Curiosity did not kill the robot: a curiosity-based learning system for a shopkeeper robot. ACM Transactions on Human–Robot Interaction (THRI), 8(3):15:1–15:24, August 2019. CODEN ???? ISSN 2573-9522 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3326462.

Dennler:2023:DMU

[DRH^{+}23] Nathaniel Dennler, Changxiao Ruan, Jessica Hadiwijoyo, Brenna Chen, Stefanos Nikolaidis, and Maja Matarić. Design metaphors for understanding user expectations of socially interactive robot embodiments. ACM Transactions on Human–Robot Interaction (THRI), 12(2):21:1–21:??, June 2023. CODEN ???? ISSN 2573-9522 (electronic). URL

https://dl.acm.org/doi/10.
1145/3550489.

Ruiz-del-Solar:2013:BGP

[dSCV^{+}13] Javier Ruiz del Solar, Mauricio Correa, Rodrigo Verschae, Fernando Bernuy, Patricio Loncomilla, Mauricio Mascaró, Romina Riquelme, and Felipe Smith. Bender: a generalpurpose social robot with human–robot interaction capabilities. Journal of Human– Robot Interaction (JHRI), 1(2): 54–75, February 2013. ISSN 2163-0364. URL https://dl.

acm.org/doi/abs/10.5898/
JHRI.1.2.Ruiz-del-Solar.

Dragan:2013:TIC

[DSL13] Anca D. Dragan, Siddhartha S.

Srinivasa, and Kenton C. T. Lee. Teleoperation with intelligent and customizable interfaces. Journal of Human– Robot Interaction (JHRI), 2(2): 33–57, June 2013. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
2.2.Dragan.

Deligianis:2017:IIB

[DSMS17] Christopher Deligianis, Christopher John Stanton, Craig McGarty, and Catherine J. Stevens. The impact of intergroup bias on trust and approach behaviour towards a humanoid robot. Journal of Human–Robot Interaction (JHRI), 6(3):4–20, December 2017. ISSN 2163-

                                        0364. URL https://dl.acm.
                                        org/doi/abs/10.5898/JHRI.
                                        6.3.Deligianis.

Dominguez:2015:CDA

[DSPM15] Cynthia Dominguez, Robert Strouse, Elizabeth Lerner Papautsky, and Brian Moon. Cognitive design of an application enabling remote bases to receive unmanned helicopter resupply. Journal of Human–Robot Interaction (JHRI), 4(2):50–60, September 2015. ISSN 2163-

                                        0364. URL https://dl.acm.
                                        org/doi/abs/10.5898/JHRI.
                                        4.2.Dominguez.

Deits:2013:CCI

[DTT^{+}13] Robin Deits, Stefanie Tellex, Pratiksha Thaker, Dimitar Simeonov, Thomas Kollar, and Nicholas Roy. Clarifying commands with informationtheoretic human–robot dialog. Journal of Human–Robot Interaction (JHRI), 2(2):58– 79, June 2013. ISSN 2163-

                                        0364. URL https://dl.acm.
                                        org/doi/abs/10.5898/JHRI.
                                        2.2.Deits.

Diel:2022:MAU

[DWM22] Alexander Diel, Sarah Weigelt, and Karl F. Macdorman. A meta-analysis of the uncanny valley’s independent and dependent variables. ACM Transactions on Human–Robot Interaction (THRI), 11(1):1:1– 1:33, March 2022. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3470742.

deWit:2023:DOE

[dWVK23] Jan de Wit, Paul Vogt, and Emiel Krahmer. The design and observed effects of robotperformed manual gestures: a systematic review. ACM Transactions on Human–Robot Interaction (THRI), 12(1):1:1– 1:??, March 2023. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3549530.

Engwall:2022:ILE

[ECL^{+}22] Olov Engwall, Ronald Cumbal, José Lopes, Mikael Ljung, and Linnea Månsson. Identification of low-engaged learners in robot-led second language conversations with adults. ACM Transactions on Human–Robot Interaction (THRI), 11(2): 18:1–18:??, June 2022. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3503799.

Edirisinghe:2023:FTS

[ESK23] Sachi Edirisinghe, Satoru Satake, and Takayuki Kanda. Field trial of a shopworker robot with friendly guidance and appropriate admonishments. ACM Transactions on Human–Robot Interaction (THRI), 12(3):34:1–34:??, September 2023. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3575805.

Ferrier-Barbut:2022:MQL

[FBGL^{+}22] Eléonore Ferrier-Barbut, Philippe Gauthier, Vanda Luengo, Geoffroy Canlorbe, and Marie-Aude Vitrani. Measuring the quality of learning in a human-robot collaboration: a study of laparoscopic surgery. ACM Transactions on Human– Robot Interaction (THRI), 11 (3):27:1–27:??, September 2022. CODEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3476414.

Figueredo:2022:PMH

[FDC^{+}22] Luis F. C. Figueredo, Rafael De Castro Aguiar, Lipeng Chen, Thomas C. Richards, Samit Chakrabarty, and Mehmet Dogar. Planning to minimize the human muscular effort during forceful human-robot collaboration. ACM Transactions on Human–Robot Interaction (THRI), 11(1):10:1– 10:27, March 2022. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3481587.

Forster:2020:TSM

[FDN20] Frank Förster, Kerstin Dautenhahn, and Chrystopher L. Nehaniv. Toward scalable measures of quality of interaction: Motor interference. ACM Transactions on Human–Robot Interaction (THRI), 9(2):8:1– 8:25, February 2020. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3344277.

Fulton:2022:RCM

[FES22] Michael Fulton, Chelsey Edge, and Junaed Sattar. Robot communication via motion: a study on modalities for robotto-human communication in the field. ACM Transactions on Human–Robot Interaction (THRI), 11(2):15:1–15:??, June 2022. CODEN ???? ISSN 2573-9522 (electronic). URL

https://dl.acm.org/doi/10.
1145/3495245.

Fossati:2023:RPW

[FGCB23] Maria R. Fossati, Giorgio Grioli, Manuel G. Catalano, and Antonio Bicchi. From robotics to prosthetics: What design and engineering can do better together. ACM Transactions on Human–Robot Interaction (THRI), 12(2):27:1– 27:??, June 2023. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3588323.

Fitzgerald:2018:HGO

[FGT18] Tesca Fitzgerald, Ashok Goel, and Andrea Thomaz. Humanguided object mapping for task transfer. ACM Transactions on Human–Robot Interaction (THRI), 7(2):17:1–17:24, October 2018. CODEN ???? ISSN 2573-9522 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3277905.

Fischer:2021:ECP

[Fis21] Kerstin Fischer. Effect confirmed, patient dead: a commentary on Hoffman & Zhao’s primer for conducting experiments in HRI. ACM Transactions on Human–Robot Interaction (THRI), 10(1):9:1–9:4, February 2021. CODEN ???? ISSN 2573-9522 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3439714.  See

[HZ21].

Fischer:2022:TAB

[Fis22] Kerstin Fischer. Tracking anthropomorphizing behavior in human-robot interaction. ACM Transactions on Human–Robot Interaction (THRI), 11(1):4:1– 4:28, March 2022. CODEN ???? ISSN 2573-9522 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3442677.

Ferland:2013:NID

[FLA^{+}13] François Ferland, Dominic Létourneau, Arnaud Aumont, Julien Frémy, Marc-Antoine Legault, Michel Lauria, and François Michaud. Natural interaction design of a humanoid robot. Journal of Human–Robot Interaction (JHRI), 1(2):118– 134, February 2013. ISSN 2163-

                                        0364. URL https://dl.acm.
                                        org/doi/abs/10.5898/JHRI.
                                        1.2.Ferland.

Fasola:2013:SAR

[FM13] Juan Fasola and Maja J. Matarić. A socially assistive robot exercise coach for the elderly. Journal of Human– Robot Interaction (JHRI), 2(2): 3–32, June 2013. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
2.2.Fasola.

Fitzsimons:2022:ESC

[FM22] Kathleen Fitzsimons and Todd D.

Murphey. Ergodic shared control: Closing the loop on pHRI based on information encoded in motion. ACM Transactions on Human–Robot Interaction (THRI), 11(4):37:1– 37:??, December 2022. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3526106.

Fontaine:2022:EHR

[FN22] Matthew C. Fontaine and Stefanos Nikolaidis. Evaluating human-robot interaction algorithms in shared autonomy via quality diversity scenario generation. ACM Transactions on Human–Robot Interaction (THRI), 11(3):25:1–25:??, September 2022. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3476412.

Fischer:2020:SMM

[FNJB20] Kerstin Fischer, Oliver Niebuhr, Lars C. Jensen, and Leon Bodenhagen. Speech melody matters — how robots profit from using charismatic speech. ACM Transactions on Human–Robot Interaction (THRI), 9(1):4:1– 4:21, January 2020. CODEN ???? ISSN 2573-9522 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3344274.

Feil-Seifer:2017:YRT

[FS17] David Feil-Seifer. Are you reviewer 2: Three ideas for better reviewing. Journal of Human– Robot Interaction (JHRI), 6(3): 1–3, December 2017. ISSN 2163-0364. URL https://dl.

                                        acm.org/doi/abs/10.5898/
                                        JHRI.6.3.Feil-Seifer.

Feil-Seifer:2021:WNI

[FSHR^{+}21] David Feil-Seifer, Kerstin S. Haring, Silvia Rossi, Alan R. Wagner, and Tom Williams. Where to next? The impact of COVID-19 on human–robot interaction research. ACM Transactions on Human–Robot Interaction (THRI), 10(1):1:1– 1:7, February 2021. CODEN ???? ISSN 2573-9522 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3405450.

Forster:2019:RLS

[FSLN19] Frank Förster, Joe Saunders, Hagen Lehmann, and Chrystopher L. Nehaniv. Robots learning to say “no”: Prohibition and rejective mechanisms in acquisition of linguistic negation. ACM Transactions on Human–Robot Interaction (THRI), 8(4):23:1–23:26, December 2019. CODEN ???? ISSN 2573-9522 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3359618.

Feil-Seifer:2012:DBC

[FSM12] David Feil-Seifer and Maja J.

Matarić. Distance-based computational models for facilitating robot interaction with children. Journal of Human– Robot Interaction (JHRI), 1(1): 55–77, July 2012. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
1.1.Feil-Seifer.

Fehlberg:2015:EPA

[FSP15] Mark A. Fehlberg, Hamidreza N.

Sani, and William R. Provancher. Enhancements to the planar active handrest. Journal of Human–Robot Interaction (JHRI), 4(3):149–169, December 2015. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
4.3.Fehlberg.

Gasteiger:2021:RDC

[GAG^{+}21] Norina Gasteiger, Ho Seok Ahn, Chiara Gasteiger, Christopher Lee, Jongyoon Lim, Christine Fok, Bruce A. Macdonald, Geon Ha Kim, and Elizabeth Broadbent. Robot-delivered cognitive stimulation games for older adults: Usability and acceptability evaluation. ACM Transactions on Human– Robot Interaction (THRI), 10 (4):33:1–33:18, December 2021. CODEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3451882.

Gasteiger:2022:PDD

[GAL^{+}22] Norina Gasteiger, Ho Seok Ahn, Christopher Lee, Jongyoon Lim, Bruce A. MacDonald, Geon Ha Kim, and Elizabeth Broadbent. Participatory design, development, and testing of assistive health robots with older adults: an international four-year project. ACM Transactions on Human–Robot Interaction (THRI), 11(4):45:1– 45:??, December 2022. CODEN ???? ISSN 2573-9522 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3533726.

Gardenfors:2019:UER

[Gär19] Peter Gärdenfors. Using event representations to generate robot semantics. ACM Transactions on Human–Robot Interaction (THRI), 8(4):21:1– 21:21, December 2019. CO-DEN ???? ISSN 2573-9522 (electronic). URL https://dl.

                                        acm.org/doi/abs/10.1145/
                                        3341167.

Gleeson:2015:TPA

[GCMC15] Brian Gleeson, Katelyn Currie, Karon MacLean, and Elizabeth Croft. Tap and push: assessing the value of direct physical control in human–robot collaborative tasks. Journal of Human– Robot Interaction (JHRI), 4 (1):95–113, July 2015. ISSN 2163-0364. URL https://dl.

                                        acm.org/doi/abs/10.5555/
                                        3109835.3109841.

Geiskkovitch:2016:PCW

[GCSY16] Denise Y. Geiskkovitch, Derek Cormier, Stela H. Seo, and James E. Young. Please continue, we need more data: an exploration of obedience to robots. Journal of Human– Robot Interaction (JHRI), 5 (1):82–99, March 2016. ISSN 2163-0364. URL https://dl.

acm.org/doi/abs/10.5555/
3109939.3109943.

Goodrich:2013:ISIa

[GE13] Michael A. Goodrich and Kerstin Severinson Eklundh. Introduction to the special issue on HRI perspectives and projects from around the globe. Journal of Human–Robot Interaction (JHRI), 1(2):1–3, February 2013. ISSN 2163-0364.

URL https://dl.acm.org/
doi/abs/10.5898/JHRI.1.2.
Goodrich.

Ghafurian:2021:SRC

[GHD21] Moojan Ghafurian, Jesse Hoey, and Kerstin Dautenhahn. Social robots for the care of persons with dementia: a systematic review. ACM Transactions on Human–Robot Interaction (THRI), 10(4):41:1– 41:31, December 2021. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3469653.

Goodrich:2013:ISIb

[Goo13] Michael A. Goodrich. Introduction to the special issue on technical and social advances in HRI: an invitational issue of JHRI. Journal of Human– Robot Interaction (JHRI), 2(2): 1–2, June 2013. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
2.2.Goodrich.

Guggenheim:2020:LGI

[GPFA20] Jacob W. Guggenheim, Federico Parietti, Tamar Flash, and H. Harry Asada. Laying the groundwork for intra-robotic-natural limb coordination: Is fully manual control viable? ACM Transactions on Human– Robot Interaction (THRI), 9(3): 18:1–18:12, July 2020. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3377329.

Gorur:2023:FFD

[GRSA23] O. Can Görür, Benjamin Rosman, Fikret Sivrikaya, and Sahin Albayrak. FABRIC: a framework for the design and evaluation of collaborative robots with extended human adaptation. ACM Transactions on Human–Robot Interaction (THRI), 12(3):38:1–38:??, September 2023. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3585276.

Glas:2013:NRS

[GSF^{+}13] Dylan F. Glas, Satoru Satake, Florent Ferreri, Takayuki Kanda, Norihiro Hagita, and Hiroshi Ishiguro. The network robot system: enabling social human–robot interaction in public spaces. Journal of Human–Robot Interaction (JHRI), 1(2):5–32, February 2013. ISSN 2163-0364.

URL https://dl.acm.org/
doi/abs/10.5555/3109688.
3109690.

Hartzog:2016:TAR

[Har16] Woodrow Hartzog. Et tu, Android?: regulating dangerous and dishonest robots. Journal of Human–Robot Interaction (JHRI), 5(3):70–81, December 2016. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
5.3.Hartzog.

Howard:2018:HHB

[HB18] Ayanna Howard and Jason Borenstein. Hacking the human bias in robotics. ACM Transactions on Human–Robot Interaction (THRI), 7(1):3:1– 3:3, May 2018. CODEN ???? ISSN 2573-9522 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3208974.

Harriott:2015:MWT

[HBAZ15] Caroline E. Harriott, Glenna L.

Buford, Julie A. Adams, and Tao Zhang. Mental workload and task performance in peer-based human–robot teams. Journal of Human– Robot Interaction (JHRI), 4(2): 61–96, September 2015. ISSN 2163-0364. URL https://dl.

acm.org/doi/abs/10.5555/
3109842.3109846.

Hayes:2018:ISI

[HCR18] Bradley Hayes, Maya Cakmak, and Stephanie Rosenthal. Introduction to the special issue on artificial intelligence and human–robot interaction. ACM Transactions on Human–Robot Interaction (THRI), 7(2):13:1–13:3, October 2018. CODEN ???? ISSN 2573-9522 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3279995.

Holmquist:2014:IJH

[HF14] Lars Erik Holmquist and Jodi Forlizzi. Introduction to Journal of Human–Robot Interaction special issue on design. Journal of Human–Robot Interaction (JHRI), 3(1):1–3, February 2014. ISSN 2163-

                                        0364. URL https://dl.acm.
                                        org/doi/abs/10.5898/JHRI.
                                        3.1.Holmquist.

Herzog:2022:IRD

[HFK^{+}22] Olivia Herzog, Niklas Forchhammer, Penny Kong, Philipp Maruhn, Henriette Cornet, and Fritz Frenkler. The influence of robot designs on human compliance and emotion: a virtual reality study in the context of future public transport. ACM Transactions on Human– Robot Interaction (THRI), 11 (2):21:1–21:??, June 2022. CO-DEN ???? ISSN 2573-9522 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3507472.

Han:2021:BFR

[HGA^{+}21] Zhao Han, Daniel Giger, Jordan Allspaw, Michael S. Lee, Henny Admoni, and Holly A. Yanco. Building the foundation of robot explanation generation using behavior trees. ACM Transactions on Human– Robot Interaction (THRI), 10 (3):26:1–26:31, July 2021. CO-DEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3457185.

Hu:2023:WCR

[HH23] Yuhan Hu and Guy Hoffman. What can a robot’s skin be? Designing texturechanging skin for human-robot social interaction. ACM Transactions on Human–Robot Interaction (THRI), 12(2):26:1– 26:??, June 2023. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3532772.

Heard:2019:DHW

[HHHA19] Jamison Heard, Rachel Heald, Caroline E. Harriott, and Julie A. Adams. A diagnostic human workload assessment algorithm for collaborative and supervisory human– robot teams. ACM Transactions on Human–Robot Interaction (THRI), 8(2):7:1–7:30, June 2019. CODEN ???? ISSN 2573-9522 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3314387.

Hoffman:2014:DRM

[HJ14] Guy Hoffman and Wendy Ju. Designing robots with movement in mind. Journal of Human–Robot Interaction (JHRI), 3(1):91–122, February 2014. ISSN 2163-

                                        0364. URL https://dl.acm.
                                        org/doi/abs/10.5898/JHRI.
                                        3.1.Hoffman.

Harrison:2023:ACP

[HJ23] Katherine Harrison and Ericka Johnson. Affective corners as a problematic for design interactions. ACM Transactions on Human–Robot Interaction (THRI), 12(4):41:1–41:??, 2023. CODEN ???? ISSN 2573-9522 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3596452.

Habibian:2022:HWI

[HJL22] Soheil Habibian, Ananth Jonnavittula, and Dylan P. Losey. Here’s what I’ve learned: Asking questions that reveal reward learning. ACM Transactions on Human–Robot Interaction (THRI), 11(4):40:1– 40:??, December 2022. CODEN ???? ISSN 2573-9522 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3526107.

Huber:2014:DAR

[HLWV14] Andreas Huber, Lara Lammer, Astrid Weiss, and Markus Vincze. Designing adaptive roles for socially assistive robots: a new method to reduce technological determinism and role stereotypes. Journal of Human–Robot Interaction (JHRI), 3(2):100– 115, July 2014. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
3.2.Huber.

Huang:2013:RRB

[HM13] Chien-Ming Huang and Bilge Mutlu. The repertoire of robot behavior: enabling robots to achieve interaction goals through social behavior. Journal of Human–Robot Interaction (JHRI), 2(2):80–102, June 2013. ISSN 2163-0364.

URL https://dl.acm.org/
doi/abs/10.5898/JHRI.2.2.
Huang.

Harutyunyan:2013:CTH

[HMGC13] Vahagn Harutyunyan, Vimitha Manohar, Issak Gezehei, and Jacob W. Crandall. Cognitive telepresence in human– robot interactions. Journal of Human–Robot Interaction (JHRI), 1(2):158–182, February 2013. ISSN 2163-0364.

URL https://dl.acm.org/
doi/abs/10.5898/JHRI.1.2.
Harutyunyan.

Han:2021:NVR

[HPY21] Zhao Han, Elizabeth Phillips, and Holly A. Yanco. The need for verbal robot explanations and how people would like a robot to explain itself. ACM Transactions on Human– Robot Interaction (THRI), 10 (4):36:1–36:42, December 2021. CODEN ???? ISSN 2573-9522 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3469652.

He:2015:IHR

[HS15] Wuwei He and Daniel Sidobre. Improving human–robot object exchange by online force classification. Journal of Human– Robot Interaction (JHRI), 4 (1):75–94, July 2015. ISSN 2163-0364. URL https://dl.

                                        acm.org/doi/abs/10.5555/
                                        3109835.3109840.

Han:2023:CMC

[HY23] Zhao Han and Holly Yanco. Communicating missing causal information to explain a robot’s past behavior. ACM Transactions on Human–Robot Interaction (THRI), 12(1):10:1– 10:??, March 2023. CODEN ???? ISSN 2573-9522 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3568024.

Hoffman:2021:PCE

[HZ21] Guy Hoffman and Xuan Zhao. A primer for conducting experiments in human–robot interaction. ACM Transactions on Human–Robot Interaction (THRI), 10(1):6:1–6:31, February 2021. CODEN ???? ISSN 2573-9522 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3412374. See comments

[Fis21, SVD21].

Irfan:2022:MMO

[IOLB22] Bahar Irfan, Michael Garcia Ortiz, Natalia Lyubova, and Tony Belpaeme. Multi-modal open world user identification. ACM Transactions on Human– Robot Interaction (THRI), 11 (1):6:1–6:50, March 2022. CO-DEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3477963.

Itoh:2015:TTH

[ITI15] Makoto Itoh, Hiroto Tanaka, and Toshiyuki Inagaki. Toward trustworthy haptic assistance system for emergency avoidance of collision with a pedestrian. Journal of Human–Robot Interaction (JHRI), 4(3):4–18, December 2015. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
4.3.Itoh.

Jain:2020:PHI

[JA20] Siddarth Jain and Brenna Argall. Probabilistic human intent recognition for shared autonomy in assistive robotics. ACM Transactions on Human– Robot Interaction (THRI), 9(1): 2:1–2:23, January 2020. CO-DEN ???? ISSN 2573-9522 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/
3359614.

Johanson:2023:EHR

[JAG^{+}23] Deborah Johanson, Ho Seok Ahn, Rishab Goswami, Kazuki Saegusa, and Elizabeth Broadbent. The effects of healthcare robot empathy statements and head nodding on trust and satisfaction: a video study. ACM Transactions on Human–Robot Interaction (THRI), 12(1):4:1– 4:??, March 2023. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3549534.

Johnson:2014:CDD

[JBF^{+}14] Matthew Johnson, Jeffrey M. Bradshaw, Paul J. Feltovich, Catholijn M. Jonker, M. Birna van Riemsdijk, and Maarten Sierhuis. Coactive design: designing support for interdependence in joint activity. Journal of Human–Robot Interaction (JHRI), 3(1):43–69, February 2014. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
3.1.Johnson.

Javed:2020:RFF

[JBJ^{+}20] Hifza Javed, Rachael Burns, Myounghoon Jeon, Ayanna M. Howard, and Chung Hyuk Park. A robotic framework to facilitate sensory experiences for children with autism spectrum disorder: a preliminary study. ACM Transactions on Human–Robot Interaction (THRI), 9(1):3:1–3:26, January 2020. CODEN ???? ISSN 2573-9522 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3359613.

Jung:2021:RAT

[JDS^{+}21] Malte F. Jung, Dominic Difranzo, Solace Shen, Brett Stoll, Houston Claure, and Austin Lawrence. Robotassisted tower construction — a method to study the impact of a robot’s allocation behavior on interpersonal dynamics and collaboration in groups. ACM Transactions on Human–Robot Interaction (THRI), 10(1):2:1– 2:23, February 2021. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3394287.

Jung:2018:RWT

[JH18] Malte Jung and Pamela Hinds.

Robots in the wild: a time for more robust theories of human–robot interaction. ACM Transactions on Human–Robot Interaction (THRI), 7(1):2:1– 2:5, May 2018. CODEN ???? ISSN 2573-9522 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3208975.

Jaeger:2016:IAT

[JL16] Christopher Brett Jaeger and Daniel T. Levin. If Asimo thinks, does Roomba feel?: the legal implications of attributing agency to technology. Journal of Human–Robot Interaction (JHRI), 5(3):3–25, December 2016. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
5.3.Jaeger.

Jung:2019:WDO

[JL19] Merel M. Jung and Geke D. S.

Ludden. What do older adults and clinicians think about traditional mobility aids and exoskeleton technology? ACM Transactions on Human–Robot Interaction (THRI), 8(2):10:1– 10:17, June 2019. CODEN ???? ISSN 2573-9522 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3311789.

Joosse:2021:MAH

[JLV^{+}21] Michiel Joosse, Manja Lohse, Niels Van Berkel, Aziez Sardar, and Vanessa Evers. Making appearances: How robots should approach people. ACM Transactions on Human–Robot Interaction (THRI), 10(1):7:1– 7:24, February 2021. CODEN ???? ISSN 2573-9522 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3385121.

Jorgensen:2023:TSS

[Jør23] Jonas Jørgensen. Towards a soft science of soft robots. A call for a place for aesthetics in soft robotics research. ACM Transactions on Human–Robot Interaction (THRI), 12(2):15:1– 15:??, June 2023. CODEN ???? ISSN 2573-9522 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3533681.

Jenkins:2017:EII

[JS17] Odest Chadwicke Jenkins and Selma Sabanović. Editorial introduction: impact, sustainability, and inclusion for JHRI. Journal of Human– Robot Interaction (JHRI), 6(1): 1–3, May 2017. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
6.1.Jenkins.

Jenkins:2018:UAT

[JS18] Odest Chadwicke Jenkins and Selma Sabanović. Understanding the ACM THRI review process. ACM Transactions on Human–Robot Interaction (THRI), 7(3):18:1–18:2, December 2018. CODEN ???? ISSN 2573-9522 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3289154.

Jenkins:2019:EIA

[JS19] Odest Chadwicke Jenkins and Selma Sabanović. Editorial introduction to ACM THRI volume 8, issue 2. ACM Transactions on Human–Robot Interaction (THRI), 8(2):6:1, June 2019. CODEN ???? ISSN 2573-9522 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3328224.

Jia:2021:TPT

[JS21] Shengxin Jia and Veronica J.

Santos. Tactile perception for teleoperated robotic exploration within granular media. ACM Transactions on Human– Robot Interaction (THRI), 10 (4):34:1–34:27, December 2021. CODEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3459996.

Jois:2021:WHW

[JW21] Himavath Jois and Alan R.

Wagner. What happens when robots punish? evaluating human task performance during robot-initiated punishment. ACM Transactions on Human– Robot Interaction (THRI), 10 (4):38:1–38:18, December 2021. CODEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3472207.

Jackson:2022:EMS

[JW22] Ryan Blake Jackson and Tom Williams. Enabling morally sensitive robotic clarification requests. ACM Transactions on Human–Robot Interaction (THRI), 11(2):17:1–17:??, June 2022. CODEN ???? ISSN 2573-9522 (electronic). URL

https://dl.acm.org/doi/10.
1145/3503795.

Kirchner:2013:RCP

[KA13] Nathan Kirchner and Alen Alempijevic. A robot centric perspective on the HRI paradigm. Journal of Human– Robot Interaction (JHRI), 1(2): 135–157, February 2013. ISSN 2163-0364. URL https://dl.

acm.org/doi/abs/10.5898/
JHRI.1.2.Kirchner.

Korn:2021:UCP

[KAG21] Oliver Korn, Neziha Akalin, and Ruben Gouveia. Understanding cultural preferences for social robots: a study in German and Arab communities. ACM Transactions on Human–Robot Interaction (THRI), 10(2):12:1–12:19, May 2021. CODEN ???? ISSN 2573-9522 (electronic). URL

https://dl.acm.org/doi/10.
1145/3439717.

Kanda:2013:ISI

[KB13] Takayuki Kanda and Tony Belpaeme. Introduction to the special issue on HRI system studies. Journal of Human–Robot Interaction (JHRI), 2(1):1–3, February 2013. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
2.1.Kanda.

Kunold:2023:ARE

[KBvdP23] Laura Kunold, Nikolai Bock, and Astrid Rosenthal von der Pütten. Not all robots are evaluated equally: The impact of morphological features on robots’ assessment through capability attributions. ACM Transactions on Human–Robot Interaction (THRI), 12(1):3:1– 3:??, March 2023. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3549532.

Kutbi:2021:USE

[KDC^{+}21] Mohammed Kutbi, Xiaoxue Du, Yizhe Chang, Bo Sun, Nikolaos Agadakos, Haoxiang Li, Gang Hua, and Philippos Mordohai. Usability studies of an egocentric visionbased robotic wheelchair. ACM Transactions on Human–Robot Interaction (THRI), 10(1):4:1– 4:23, February 2021. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3399434.

Krening:2018:IAE

[KF18] Samantha Krening and Karen M. Feigh. Interaction algorithm effect on human experience with reinforcement learning. ACM Transactions on Human–Robot Interaction (THRI), 7(2):16:1– 16:22, October 2018. CODEN ???? ISSN 2573-9522 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3277904.

Kim:2021:GLG

[KF21] Lawrence H. Kim and Sean Follmer. Generating legible and glanceable swarm robot motion through trajectory, collective behavior, and pre-attentive processing features. ACM Transactions on Human–Robot Interaction (THRI), 10(3):21:1– 21:25, July 2021. CODEN ???? ISSN 2573-9522 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3442681.

Kiesler:2018:SHR

[KG18] Sara Kiesler and Michael A. Goodrich. The science of human–robot interaction. ACM Transactions on Human–Robot Interaction (THRI), 7(1):9:1– 9:3, May 2018. CODEN ???? ISSN 2573-9522 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3209701.

Kaptein:2022:CBR

[KKC^{+}22] Frank Kaptein, Bernd Kiefer, Antoine Cully, Oya Celiktutan, Bert Bierman, Rifca Rijgersberg-peters, Joost Broekens, Willeke Van Vught, Michael Van Bekkum, Yiannis Demiris, and Mark A. Neerincx. A cloud-based robot system for long-term interaction: Principles, implementation, lessons learned. ACM Transactions on Human–Robot Interaction (THRI), 11(1):8:1–8:27, March 2022. CODEN ???? ISSN 2573-9522 (electronic). URL

https://dl.acm.org/doi/10.
1145/3481585.

Karreman:2019:BRD

[KLE19] Daphne E. Karreman, Geke D. S. Ludden, and Vanessa Evers. Beyond R2D2: Designing multimodal interaction behavior for robot-specific morphology. ACM Transactions on Human–Robot Interaction (THRI), 8(3):18:1–18:32, August 2019. CODEN ???? ISSN 2573-9522 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3331144.

Krausman:2022:TMH

[KNF^{+}22] Andrea Krausman, Catherine Neubauer, Daniel Forster, Shan Lakhmani, Anthony L. Baker, Sean M. Fitzhugh, Gregory Gremillion, Julia L. Wright, Jason S. Metcalfe, and Kristin E. Schaefer. Trust measurement in human-autonomy teams: Development of a conceptual toolkit. ACM Transactions on Human–Robot Interaction (THRI), 11(3):33:1–33:??, September 2022. CODEN ???? ISSN 2573-9522 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3530874.

Kim:2012:BRG

[KPSS12] Elizabeth S. Kim, Rhea Paul, Frederick Shic, and Brian Scassellati. Bridging the research gap: making HRI useful to individuals with autism. Journal of Human–Robot Interaction (JHRI), 1(1):26–54, July 2012. ISSN 2163-0364.

                                        URL https://dl.acm.org/
                                        doi/abs/10.5898/JHRI.1.1.
                                        Kim.

Kolling:2013:HSI

[KSNL13] Andreas Kolling, Katia Sycara, Steven Nunnally, and Michael Lewis. Human–swarm interaction: an experimental study of two types of interaction with foraging swarms. Journal of Human–Robot Interaction (JHRI), 2(2):103–129, June 2013. ISSN 2163-0364.

                                        URL https://dl.acm.org/
                                        doi/abs/10.5898/JHRI.2.2.
                                        Kolling.

Kondo:2013:GCA

[KTTO13] Yutaka Kondo, Kentaro Takemura, Jun Takamatsu, and Tsukasa Ogasawara. A gesturecentric Android system for multi-party human–robot interaction. Journal of Human– Robot Interaction (JHRI), 2(1): 133–151, February 2013. ISSN 2163-0364. URL https://dl.

                                        acm.org/doi/abs/10.5898/
                                        JHRI.2.1.Kondo.

Lemasurier:2021:MER

[LBA^{+}21] Gregory Lemasurier, Gal Bejerano, Victoria Albanese, Jenna Parrillo, Holly A. Yanco, Nicholas Amerson, Rebecca Hetrick, and Elizabeth Phillips. Methods for expressing robot intent for human-robot collaboration in shared workspaces. ACM Transactions on Human– Robot Interaction (THRI), 10 (4):40:1–40:27, December 2021. CODEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3472223.

Li:2019:CDN

[LCM^{+}19] Jamy Li, Andrea Cuadra, Brian Mok, Byron Reeves, Jofish Kaye, and Wendy Ju. Communicating dominance in a nonanthropomorphic robot using locomotion. ACM Transactions on Human–Robot Interaction (THRI), 8(1):4:1–4:14, March 2019. CODEN ???? ISSN 2573-9522 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3310357.

Law:2022:EAR

[LCTRS22] Theresa Law, Meia Chita-Tegmark, Nicholas Rabb, and Matthias Scheutz. Examining attachment to robots: Benefits, challenges, and alternatives. ACM Transactions on Human– Robot Interaction (THRI), 11 (4):36:1–36:??, December 2022. CODEN ???? ISSN 2573-9522 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3526105.

Levin:2013:CDM

[LHP^{+}13] Daniel T. Levin, Caroline Harriott, Natalie A. Paul, Tao Zhang, and Julie A. Adams. Cognitive dissonance as a measure of reactions to human– robot interaction. Journal of Human–Robot Interaction (JHRI), 2(3):3–17, September 2013. ISSN 2163-0364.

                                        URL https://dl.acm.org/
                                        doi/abs/10.5898/JHRI.2.3.
                                        Levin.

Li:2017:TMB

[LJR17] Jamy Jue Li, Wendy Ju, and Byron Reeves. Touching a mechanical body: tactile contact with body parts of a humanoid robot is physiologically arousing. Journal of Human–Robot Interaction (JHRI), 6(3):118–130, December 2017. ISSN 2163-

                                        0364. URL https://dl.acm.
                                        org/doi/abs/10.5898/JHRI.
                                        6.3.Li.

Ljungblad:2023:ADF

[Lju23]  Sara Ljungblad.  Applying

“Designerly Framing” to understand assisted feeding as social aesthetic bodily experiences. ACM Transactions on Human–Robot Interaction (THRI), 12(2):25:1–25:??, June 2023. CODEN ???? ISSN 2573-9522 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3583742.

Lin:2022:IEE

[LKL22] Tsung-Chi Lin, Achyuthan Unni Krishnan, and Zhi Li. Intuitive, efficient and ergonomic Telenursing robot interfaces: Design evaluation and evolution. ACM Transactions on Human– Robot Interaction (THRI), 11 (3):23:1–23:??, September 2022. CODEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3526108.

Lin:2023:PMC

[LKL23] Tsung-Chi Lin, Achyuthan Unni Krishnan, and Zhi Li. Perceptionmotion coupling in active telepresence: Human behavior and teleoperation interface design. ACM Transactions on Human– Robot Interaction (THRI), 12 (3):31:1–31:??, September 2023. CODEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3571599.

Leichtmann:2022:PSH

[LLBN22] Benedikt Leichtmann, Albrecht Lottermoser, Julia Berger, and Verena Nitsch. Personal space in human-robot interaction at work: Effect of room size and working memory load. ACM Transactions on Human– Robot Interaction (THRI), 11 (4):44:1–44:??, December 2022. CODEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3536167.

Langer:2021:ERS

[LLT21] Allison Langer and Shelly Levy-Tzedek. Emerging roles for social robots in rehabilitation: Current directions. ACM Transactions on Human– Robot Interaction (THRI), 10 (4):30:1–30:4, December 2021. CODEN ???? ISSN 2573-9522 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3462256.

Lazzeri:2014:DTM

[LMD14] Nicole Lazzeri, Daniele Mazzei, and Danilo De Rossi. Development and testing of a multimodal acquisition platform for human–robot interaction affective studies. Journal of Human–Robot Interaction (JHRI), 3(2):1–24, July 2014. ISSN 2163-0364. URL https:

                                        //dl.acm.org/doi/abs/10.
                                        5898/JHRI.3.2.Lazzeri.

Liu:2022:CHR

[LNG22] Ruisen Liu, Manisha Natarajan, and Matthew C. Gombolay. Coordinating humanrobot teams with dynamic and stochastic task proficiencies. ACM Transactions on Human– Robot Interaction (THRI), 11 (1):5:1–5:42, March 2022. CO-DEN ???? ISSN 2573-9522 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3477391.

Ligthart:2023:ITT

[LNH23] Mike E. U. Ligthart, Mark A. Neerincx, and Koen V. Hindriks. It takes two: Using cocreation to facilitate child-robot co-regulation. ACM Transactions on Human–Robot Interaction (THRI), 12(4):42:1–42:??, 2023. CODEN ???? ISSN 2573-9522 (electronic). URL

https://dl.acm.org/doi/10.
1145/3593812.

Losey:2020:LCR

[LO20] Dylan P. Losey and Marcia K.

O’Malley. Learning the correct robot trajectory in realtime from physical human interactions. ACM Transactions on Human–Robot Interaction (THRI), 9(1):1:1–1:19, January 2020. CODEN ???? ISSN 2573-9522 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3354139.

Liberman-Pincu:2023:JSA

[LPPOG23] Ela Liberman-Pincu, Yisrael Parmet, and Tal Oron-Gilad. Judging a socially assistive robot by its cover: The effect of body structure, outline, and color on users’ perception. ACM Transactions on Human–Robot Interaction (THRI), 12(2):23:1–23:??, June 2023. CODEN ???? ISSN 2573-9522 (electronic). URL

https://dl.acm.org/doi/10.
1145/3571717.

Lee:2018:RAR

[LR18] Hee Rin Lee and Laurel D.

Riek. Reframing assistive robots to promote successful aging. ACM Transactions on Human–Robot Interaction (THRI), 7(1):11:1–11:23, May 2018. CODEN ???? ISSN 2573-9522 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3203303.

Lee:2023:DRA

[LR23] Hee Rin Lee and Laurel Riek. Designing robots for aging: Wisdom as a critical lens. ACM Transactions on Human–Robot Interaction (THRI), 12(1):2:1– 2:??, March 2023. CODEN ???? ISSN 2573-9522 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3549531.

Leyzberg:2018:EPL

[LRS18] Daniel Leyzberg, Aditi Ramachandran, and Brian Scassellati. The effect of personalization in longer-term robot tutoring. ACM Transactions on Human–Robot Interaction (THRI), 7(3):19:1–19:19, December 2018. CODEN ???? ISSN 2573-9522 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3283453.

Luria:2020:DCE

[LSB^{+}20] Michal Luria, Ophir Sheriff, Marian Boo, Jodi Forlizzi, and Amit Zoran. Destruction, catharsis, and emotional release in human–robot interaction. ACM Transactions on Human– Robot Interaction (THRI), 9 (4):22:1–22:19, October 2020. CODEN ???? ISSN 2573-9522 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3385007.

Lagerstedt:2023:MRM

[LT23] Erik Lagerstedt and Serge Thill. Multiple roles of multimodality among interacting agents. ACM Transactions on Human–Robot Interaction (THRI), 12(2):17:1–17:??, June 2023. CODEN ???? ISSN 2573-9522 (electronic). URL

https://dl.acm.org/doi/10.
1145/3549955.

Levillain:2017:BOR

[LZ17] Florent Levillain and Elisabetta Zibetti. Behavioral objects: the rise of the evocative machines. Journal of Human– Robot Interaction (JHRI), 6(1): 4–24, May 2017. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
6.1.Levillain.

Mulder:2015:ISI

[MAC15] Mark Mulder, David A. Abbink, and Tom Carlson. Introduction to the special issue on shared control: applications. Journal of Human–Robot Interaction (JHRI), 4(3):1–3, December 2015. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
4.3.Mulder.

Mavrogiannis:2022:SMD

[MAOTK22] Christoforos Mavrogiannis, Patrícia Alves-Oliveira, Wil Thomason, and Ross A. Knepper. Social momentum: Design and evaluation of a framework for socially competent robot navigation. ACM Transactions on Human–Robot Interaction (THRI), 11(2):14:1–14:??, June 2022. CODEN ???? ISSN 2573-9522 (electronic). URL

https://dl.acm.org/doi/10.
1145/3495244.

Mataric:2018:RBT

[Mat18] Maja Matarić. On relevance: Balancing theory and practice in HRI. ACM Transactions on Human–Robot Interaction (THRI), 7(1):8:1–8:2, May 2018. CODEN ???? ISSN 2573-9522 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3209770.

Mast:2012:UCD

[MBK^{+}12] Marcus Mast, Michael Burmester, Katja Krüger, Sascha Fatikow, Georg Arbeiter, Birgit Graf, Gernot Kronreif, Lucia Pigini, David Facal, and Renxi Qiu. User-centered design of a dynamic-autonomy remote interaction concept for manipulation-capable robots to assist elderly people in the home. Journal of Human– Robot Interaction (JHRI), 1(1): 96–118, July 2012. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
1.1.Mast.

Montebelli:2017:RHE

[MBLD17] Alberto Montebelli, Erik Billing, Jessica Lindblom, and Giulia Messina Dahlberg. Reframing HRI education: a dialogic reformulation of HRI education to promote diverse thinking and scientific progress.

Journal of Human–Robot Interaction (JHRI), 6(2):3–26, September 2017. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
6.2.Montebelli.

Mavrogiannis:2023:CCS

[MBW^{+}23] Christoforos Mavrogiannis, Francesca Baldini, Allan Wang, Dapeng Zhao, Pete Trautman, Aaron Steinfeld, and Jean Oh. Core challenges of social robot navigation: a survey. ACM Transactions on Human–Robot Interaction (THRI), 12(3):36:1– 36:??, September 2023. CO-DEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3583741.

Marvel:2022:ISI

[MBZ^{+}22] Jeremy A. Marvel, Shelly Bagchi, Megan Zimmerman, Murat Aksu, Brian Antonishek, Yue Wang, Ross Mead, Terry Fong, and Heni Ben Amor. Introduction to the special issue on test methods for human-robot teaming performance evaluations. ACM Transactions on Human– Robot Interaction (THRI), 11 (3):22:1–22:??, September 2022. CODEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3544303.

Marvel:2020:TEI

[MBZA20] Jeremy A. Marvel, Shelly Bagchi, Megan Zimmerman, and Brian Antonishek. Towards effective interface designs for collaborative HRI in manufacturing: Metrics and measures. ACM Transactions on Human– Robot Interaction (THRI), 9 (4):25:1–25:55, October 2020. CODEN ???? ISSN 2573-9522 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3385009.

Munoz:2021:PRL

[MD21] John Edison Muñoz and Kerstin Dautenhahn. Robo Ludens: a game design taxonomy for multiplayer games using socially interactive robots. ACM Transactions on Human– Robot Interaction (THRI), 10 (4):32:1–32:28, December 2021. CODEN ???? ISSN 2573-9522 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3451343.

Memar:2020:OAH

[ME20] Amirhossein H. Memar and Ehsan T. Esfahani. Objective assessment of human workload in physical human– robot cooperation using brain monitoring. ACM Transactions on Human–Robot Interaction (THRI), 9(2):13:1–13:21, February 2020. CODEN ???? ISSN 2573-9522 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3368854.

MacLean:2015:IJH

[MF15] Karon E. MacLean and Antonio Frisoli. Introduction to Journal of Human–Robot Interaction: special issue on haptics in HRI: cooperation and communication. Journal of Human– Robot Interaction (JHRI), 4(1): 1–2, July 2015. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
4.1.MacLean.

Matsuda:2016:EBM

[MHI16] Goh Matsuda, Kazuo Hiraki, and Hiroshi Ishiguro. EEGbased mu rhythm suppression to measure the effects of appearance and motion on perceived human likeness of a robot. Journal of Human– Robot Interaction (JHRI), 5 (1):68–81, March 2016. ISSN 2163-0364. URL https://dl.

acm.org/doi/abs/10.5555/
3109939.3109942.

Moon:2021:DHG

[MHV^{+}21] Ajung Moon, Maneezhay Hashmi, H. F. Machiel Van Der Loos, Elizabeth A. Croft, and Aude Billard. Design of hesitation gestures for nonverbal humanrobot negotiation of conflicts. ACM Transactions on Human– Robot Interaction (THRI), 10 (3):24:1–24:25, July 2021. CO-DEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3418302.

Ma:2022:MHR

[MIFP22] Lanssie Mingyue Ma, Martijn Ijtsma, Karen M. Feigh, and Amy R. Pritchett. Metrics for human-robot team design: a teamwork perspective on evaluation of humanrobot teams. ACM Transactions on Human–Robot Interaction (THRI), 11(3):30:1–30:??, September 2022. CODEN ???? ISSN 2573-9522 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3522581.

Miller:2013:FSC

[Mil13] Christopher A. Miller. Frameworks for supervisory control: characterizing relationships with uninhabited vehicles. Journal of Human–Robot Interaction (JHRI), 1(2):183– 200, February 2013. ISSN 2163-

                                        0364. URL https://dl.acm.
                                        org/doi/abs/10.5898/JHRI.
                                        1.2.Miller.

Mikalonyte:2022:CAI

[MK22] Elze Sigute Mikalonyte and Markus Kneer. Can artificial intelligence make art?: Folk intuitions as to whether AI-driven robots can be viewed as artists and produce art. ACM Transactions on Human–Robot Interaction (THRI), 11(4):43:1– 43:??, December 2022. CODEN ???? ISSN 2573-9522 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3530875.

Morales:2014:WTS

[MKH14] Yoichi Morales, Takayuki Kanda, and Norihiro Hagita. Walking together: side-by-side walking model for an interacting robot. Journal of Human– Robot Interaction (JHRI), 3(2): 50–73, July 2014. ISSN 2163-

                                        0364. URL https://dl.acm.

org/doi/abs/10.5898/JHRI.
3.2.Morales.

Meng:2021:LEI

[MLF^{+}21] Lingheng Meng, Daiwei Lin, Adam Francey, Rob Gorbet, Philip Beesley, and Dana Kulić. Learning to engage with interactive systems: a field study on deep reinforcement learning in a public museum. ACM Transactions on Human–Robot Interaction (THRI), 10(1):5:1– 5:29, February 2021. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3408876.

Mead:2016:RNT

[MM16] Ross Mead and Maja J.

Matarić. Robots have needs too: how and why people adapt their proxemic behavior to improve robot social signal understanding. Journal of Human–Robot Interaction (JHRI), 5(2):48–68, September 2016. ISSN 2163-0364.

URL https://dl.acm.org/
doi/abs/10.5898/JHRI.5.2.
Mead.

McColl:2013:MTS

[MN13] Derek McColl and Goldie Nejat. Meal-time with a socially assistive robot and older adults at a long-term care facility. Journal of Human–Robot Interaction (JHRI), 2(1):152–171, February 2013. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
2.1.McColl.

Moro:2018:LPS

[MNM18] Christina Moro, Goldie Nejat, and Alex Mihailidis. Learning and personalizing socially assistive robot behaviors to aid with activities of daily living. ACM Transactions on Human–Robot Interaction (THRI), 7(2):15:1–15:25, October 2018. CODEN ???? ISSN 2573-9522 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3277903.

Moon:2013:DIH

[MPCV13] AJung Moon, Chris A. C. Parker, Elizabeth A. Croft, and H. F. Machiel Van der Loos. Design and impact of hesitation gestures during human–robot resource conflicts. Journal of Human–Robot Interaction (JHRI), 2(3):18–40, September 2013. ISSN 2163-

                                        0364. URL https://dl.acm.
                                        org/doi/abs/10.5898/JHRI.
                                        2.3.Moon.

Meissner:2021:FFU

[MTCKS21] Antonia Meissner, Angelika Trübswetter, Antonia S. Conti-Kufner, and Jonas Schmidtler. Friend or foe? Understanding assembly workers’ acceptance of human–robot collaboration. ACM Transactions on Human–Robot Interaction (THRI), 10(1):3:1–3:30, February 2021. CODEN ???? ISSN 2573-9522 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3399433.

Moualeu:2015:MOE

[MU15] Antonio Moualeu and Jun Ueda. A model for operator endpoint stiffness prediction during physical human– robot interaction. Journal of Human–Robot Interaction (JHRI), 4(3):170–193, December 2015. ISSN 2163-0364.

URL https://dl.acm.org/
doi/abs/10.5898/JHRI.4.3.
Moualeu.

Mollen:2023:BCR

[MvdPD23] Joost Mollen, Peter van der Putten, and Kate Darling. Bonding with a couchsurfing robot: The impact of common locus on human-robot bonding in-the-wild. ACM Transactions on Human–Robot Interaction (THRI), 12(1):8:1–8:??, March 2023. CODEN ???? ISSN 2573-9522 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563702.

Marrella:2022:MDE

[MWIN22] Andrea Marrella, Lun Wang, Luca Iocchi, and Daniele Nardi. A methodology to design and evaluate HRI teaming tasks in robotic competitions. ACM Transactions on Human–Robot Interaction (THRI), 11(3): 34:1–34:??, September 2022. CODEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3528415.

Mubin:2019:RPS

[MWJO19] Omar Mubin, Kewal Wadibhasme, Philipp Jordan, and Mohammad Obaid. Reflecting on the presence of science fiction robots in computing literature. ACM Transactions on Human–Robot Interaction (THRI), 8(1):5:1–5:25, March 2019. CODEN ???? ISSN 2573-9522 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3303706.

Matsumoto:2022:FEP

[MWR22] Sachiko Matsumoto, Auriel Washburn, and Laurel D. Riek. A framework to explore proximate human-robot coordination. ACM Transactions on Human–Robot Interaction (THRI), 11(3):32:1–32:??, September 2022. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3526101.

Norton:2022:MRP

[NAC^{+}22] Adam Norton, Henny Admoni, Jacob Crandall, Tesca Fitzgerald, Alvika Gautam, Michael Goodrich, Amy Saretsky, Matthias Scheutz, Reid Simmons, Aaron Steinfeld, and Holly Yanco. Metrics for robot proficiency self-assessment and communication of proficiency in human-robot teams. ACM Transactions on Human–Robot Interaction (THRI), 11(3): 29:1–29:??, September 2022. CODEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3522579.

Nostadt:2020:EPT

[NACB20] Nicolas Nostadt, David A.

Abbink, Oliver Christ, and Philipp Beckerle. Embodiment, presence, and their intersections: Teleoperation and beyond. ACM Transactions on Human–Robot Interaction (THRI), 9(4):28:1–28:19, October 2020. CODEN ???? ISSN 2573-9522 (electronic). URL

https://dl.acm.org/doi/10.
1145/3389210.

Nikolaidis:2018:PVC

[NKFS18] Stefanos Nikolaidis, Minae Kwon, Jodi Forlizzi, and Siddhartha Srinivasa. Planning with verbal communication for human–robot collaboration. ACM Transactions on Human–Robot Interaction (THRI), 7(3):22:1–22:21, December 2018. CODEN ???? ISSN 2573-9522 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3203305.

Noguchi:2020:PTS

[NKT20] Yohei Noguchi, Hiroko Kamide, and Fumihide Tanaka. Personality traits for a social mediator robot encouraging elderly selfdisclosure on loss experiences. ACM Transactions on Human– Robot Interaction (THRI), 9(3): 17:1–17:24, July 2020. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3377342.

Novoa:2021:ASR

[NMW^{+}21] José Novoa, Rodrigo Mahu, Jorge Wuth, Juan Pablo Escudero, Josué Fredes, and Néstor Becerra Yoma. Automatic speech recognition for indoor HRI scenarios. ACM Transactions on Human–Robot Interaction (THRI), 10(2): 17:1–17:30, May 2021. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3442629.

Nielsen:2023:UUG

[NSHK23] Sara Nielsen, Mikael B. Skov, Karl Damkjær Hansen, and Aleksandra Kaszowska. Using user-generated YouTube videos to understand unguided interactions with robots in public places. ACM Transactions on Human–Robot Interaction (THRI), 12(1):5:1–5:??, March 2023. CODEN ???? ISSN 2573-9522 (electronic). URL

https://dl.acm.org/doi/10.
1145/3550280.

Nishimura:2015:HSC

[NWS15] Ryota Nishimura, Takahiro Wada, and Seiji Sugiyama. Haptic shared control in steering operation based on cooperative status between a driver and a driver assistance system. Journal of Human–Robot Interaction (JHRI), 4(3):19– 37, December 2015. ISSN 2163-0364. URL https://dl.

acm.org/doi/abs/10.5555/
3109848.3109851.

Onnasch:2022:IAR

[OH22] Linda Onnasch and Clara Laudine Hildebrandt. Impact of anthropomorphic robot design on trust and attention in industrial human-robot interaction. ACM Transactions on Human–Robot Interaction (THRI), 11(1):2:1– 2:24, March 2022. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3472224.

Okamura:2018:HDH

[Oka18] Allison M. Okamura. Haptic dimensions of human–robot interaction. ACM Transactions on Human–Robot Interaction (THRI), 7(1):6:1–6:3, May 2018. CODEN ???? ISSN 2573-9522 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3209768.

Oh:2014:MME

[OP14] Chang Geun Oh and Jaeheung Park. From mechanical metamorphosis to empathic interaction: a historical overview of robotic creatures. Journal of Human–Robot Interaction (JHRI), 3(1):4–19, February 2014. ISSN 2163-0364.

URL https://dl.acm.org/
doi/abs/10.5898/JHRI.3.1.
Oh.

Olatunji:2022:LAM

[OPK^{+}22] Samuel A. Olatunji, Andre Potenza, Andrey Kiselev, Tal Oron-Gilad, Amy Loutfi, and Yael Edan. Levels of automation for a mobile robot teleoperated by a caregiver. ACM Transactions on Human– Robot Interaction (THRI), 11 (2):20:1–20:??, June 2022. CO-DEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3507471.

Oguz:2019:OHH

[ORPW19] Ozgur S. Oguz, Wolfgang Rampeltshammer, Sebastian Paillan, and Dirk Wollherr. An ontology for human-human interactions and learning interaction behavior policies. ACM Transactions on Human–Robot Interaction (THRI), 8(3):16:1– 16:26, August 2019. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3326539.

Paterson:2023:IRT

[Pat23] Mark Paterson. Inviting robot touch (by design). ACM Transactions on Human–Robot Interaction (THRI), 12(2):16:1– 16:??, June 2023. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3549533.

Putten:2018:DVS

[PB18] Astrid Rosenthal-Von Der Pütten and Nikolai Bock. Development and validation of the self-efficacy in human– robot-interaction scale (SE-HRI). ACM Transactions on Human–Robot Interaction (THRI), 7(3):21:1–21:30, December 2018. CODEN ???? ISSN 2573-9522 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3139352.

Parasuraman:2017:NUT

[PCB^{+}17] Ramviyas Parasuraman, Sergio Caccamo, Fredrik Båberg, Petter Ögren, and Mark Neerincx. A new UGV teleoperation interface for improved awareness of network connectivity and physical surroundings. Journal of Human–Robot Interaction (JHRI), 6(3):48–70, December 2017. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
6.3.Parasuraman.

Perelman:2020:WDY

[PES20] Brandon S. Perelman, Arthur W.

Evans III, and Kristin E. Schaefer. Where do you think you’re going?: Characterizing spatial mental models from planned routes. ACM Transactions on Human–Robot Interaction (THRI), 9(4):23:1–23:55, October 2020. CODEN ???? ISSN 2573-9522 (electronic). URL

https://dl.acm.org/doi/10.
1145/3385008.

Paez:2022:HRS

[PG22] John Páez and Enrique González.

Human-robot scaffolding: an architecture to Foster problemsolving skills. ACM Transactions on Human–Robot Interaction (THRI), 11(3):24:1–24:??, September 2022. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3526109.

Paterson:2023:ISI

[PHZ23] Mark Paterson, Guy Hoffman, and Caroline Yan Zheng. Introduction to the special issue on “Designing the Robot Body: Critical Perspectives on Affective Embodied Interaction”. ACM Transactions on Human–Robot Interaction (THRI), 12(2):14:1–14:??, June 2023. CODEN ???? ISSN 2573-9522 (electronic). URL

https://dl.acm.org/doi/10.
1145/3594713.

Phillips:2017:SSA

[PJ17] Elizabeth Kathleen Phillips and Florian G. Jentsch. Supporting situation awareness through robot-to-human information exchanges under conditions of visuospatial perspective taking. Journal of Human–Robot Interaction (JHRI), 6(3):92–117, December 2017. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
6.3.Phillips.

Pou-Prom:2020:CRO

[PPRR20] Chloé Pou-Prom, Stefania Raimondo, and Frank Rudzicz. A conversational robot for older adults with Alzheimer’s disease. ACM Transactions on Human– Robot Interaction (THRI), 9(3): 21:1–21:25, July 2020. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3380785.

Phillips:2016:HAT

[PSB^{+}16] Elizabeth Phillips, Kristin E. Schaefer, Deborah R. Billings, Florian Jentsch, and Peter A.

Hancock. Human-animal teams as an analog for future human– robot teams: influencing design and fostering trust. Journal of Human–Robot Interaction (JHRI), 5(1):100–125, March 2016. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
5.1.Phillips.

Pena:2020:HPS

[PT20] Denis Peña and Fumihide Tanaka. Human perception of social robot’s emotional states via facial and thermal expressions. ACM Transactions on Human–Robot Interaction (THRI), 9(4):26:1–26:19, October 2020. CODEN ???? ISSN 2573-9522 (electronic). URL

https://dl.acm.org/doi/10.
1145/3388469.

Randall:2020:SRA

[Ran20] Natasha Randall. A survey of Robot-Assisted Language Learning (RALL). ACM Transactions on Human–Robot Interaction (THRI), 9(1):7:1–7:36, January 2020. CODEN ???? ISSN 2573-9522 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3345506.

Rakhsha:2015:APC

[RC15] Ramtin Rakhsha and Daniela Constantinescu. Averageposition coordination for distributed multi-user networked haptic cooperation. Journal of Human–Robot Interaction (JHRI), 4(3):62–75, December 2015. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
4.3.Rakhsha.

Roundtree:2022:TIH

[RCL^{+}22] Karina A. Roundtree, Jason R. Cody, Jennifer Leaf, H. Onan Demirel, and Julie A. Adams. Transparency’s influence on human-collective interactions. ACM Transactions on Human– Robot Interaction (THRI), 11 (2):19:1–19:??, June 2022. CO-DEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3507470.

Riley:2023:FHR

[RF23] Danny G. Riley and Eric W. Frew. Fielded human-robot interaction for a heterogeneous team in the DARPA subterranean challenge. ACM Transactions on Human–Robot Interaction (THRI), 12(3):40:1– 40:??, September 2023. CO-DEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3588325.

Reveleau:2015:VRI

[RFL^{+}15] Aurélien Reveleau, François Ferland, Mathieu Labbé, Dominic Létourneau, and François Michaud. Visual representation of interaction force and sound source in a teleoperation user interface for a mobile robot. Journal of Human–Robot Interaction (JHRI), 4(2):1–23, September 2015. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
4.2.Reveleau.

Rousseau:2013:SIM

[RFLM13] Vincent Rousseau, François Ferland, Dominic Létourneau, and François Michaud. Sorry to interrupt, but may I have your attention?: preliminary design and evaluation of autonomous engagement in HRI. Journal of Human–Robot Interaction (JHRI), 2(3):41–61, September 2013. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
2.3.Rousseau.

Riek:2012:WOS

[Rie12] Laurel D. Riek. Wizard of Oz studies in HRI: a systematic review and new reporting guidelines. Journal of Human–Robot Interaction (JHRI), 1(1):119– 136, July 2012. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
1.1.Riek.

Rueben:2021:MMM

[RKD^{+}21] Matthew Rueben, Jeffrey Klow, Madelyn Duer, Eric Zimmerman, Jennifer Piacentini, Madison Browning, Frank J. Bernieri, Cindy M. Grimm, and William D. Smart. Mental models of a mobile shoe rack: Exploratory findings from a long-term in-the-wild study. ACM Transactions on Human– Robot Interaction (THRI), 10 (2):16:1–16:36, May 2021. CO-DEN ???? ISSN 2573-9522 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3442620.

Roesler:2023:EMS

[RMO23] Eileen Roesler, Dietrich Manzey, and Linda Onnasch. Embodiment matters in social HRI research: Effectiveness of anthropomorphism on subjective and objective outcomes. ACM Transactions on Human–Robot Interaction (THRI), 12(1):7:1– 7:??, March 2023. CODEN ???? ISSN 2573-9522 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3555812.

Reinhardt:2021:BER

[RPB21] Jakob Reinhardt, Lorenz Prasch, and Klaus Bengler. Backoff: Evaluation of robot motion strategies to facilitate humanrobot spatial interaction. ACM Transactions on Human–Robot Interaction (THRI), 10(3): 23:1–23:25, July 2021. CODEN ???? ISSN 2573-9522 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3418303.

Roche:2022:SKN

[RSB22] Lucas Roche and Ludovic Saint-Bauzel. Study of kinesthetic negotiation ability in lightweight comanipulative decision-making tasks: Design and study of a virtual partner based on human– human interaction observation. ACM Transactions on Human– Robot Interaction (THRI), 11 (2):12:1–12:??, June 2022. CO-DEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3485753.

Robinson:2023:RVH

[RTC^{+}23] Nicole Robinson, Brendan Tidd, Dylan Campbell, Dana Kulić, and Peter Corke. Robotic vision for human-robot interaction and collaboration: a survey and systematic review. ACM Transactions on Human– Robot Interaction (THRI), 12 (1):12:1–12:??, March 2023. CODEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3570731.

Rudaz:2023:IOA

[RTSL23] Damien Rudaz, Karen Tatarian, Rebecca Stower, and Christian Licoppe. From inanimate object to agent: Impact of pre-beginnings on the emergence of greetings with a robot. ACM Transactions on Human– Robot Interaction (THRI), 12 (3):29:1–29:??, September 2023. CODEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3575806.

Rosenthal:2022:IRD

[RVC22] Stephanie Rosenthal, Peerat Vichivanives, and Elizabeth Carter. The impact of route descriptions on human expectations for robot navigation. ACM Transactions on Human– Robot Interaction (THRI), 11 (4):35:1–35:??, December 2022. CODEN ???? ISSN 2573-9522 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3526104.

Sklar:2015:ABD

[SA15] Elizabeth I. Sklar and M. Q. Azhar. Argumentation-based dialogue games for shared control in human–robot systems. Journal of Human–Robot Interaction (JHRI), 4(3):120– 148, December 2015. ISSN 2163-0364. URL https://dl.

                                        acm.org/doi/abs/10.5555/
                                        3109848.3109856.

Sarathy:2019:WEN

[SAS19] Vasanth Sarathy, Thomas Arnold, and Matthias Scheutz. When exceptions are the norm: Exploring the role of consent in HRI. ACM Transactions on Human–Robot Interaction (THRI), 8(3):14:1–14:21, August 2019. CODEN ???? ISSN 2573-9522 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3341166.

Sabanovic:2017:ISI

[SBB17] Selma Sabanović, Carlotta A. Berry, and Cindy L. Bethel. Introduction to the special issue on HRI education. Journal of Human–Robot Interaction (JHRI), 6(2):1–2, September 2017. ISSN 2163-0364.

                                        URL https://dl.acm.org/
                                        doi/abs/10.5555/3183905.
                                        3183906.

Sanchez:2017:IDR

[SCG17] Flor Ángela Bravo Sánchez, Alejandra María González Correal, and Enrique González Guerrero. Interactive drama with robots for teaching nontechnical subjects. Journal of Human–Robot Interaction (JHRI), 6(2):48–69, September 2017. ISSN 2163-0364.

URL https://dl.acm.org/
doi/abs/10.5898/JHRI.6.2.
Bravo.

Scherer:2014:MMM

[Sch14] Derek Scherer. Movie magic makes better social robots: the overlap of special effects and character robot engineering. Journal of Human–Robot Interaction (JHRI), 3(1):123–141, February 2014. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
3.1.Scherer.

Schneiders:2022:NDI

[SCK^{+}22] Eike Schneiders, EunJeong Cheon, Jesper Kjeldskov, Matthias Rehm, and Mikael B. Skov. Non-dyadic interaction: a literature review of 15 years of human-robot interaction conference publications. ACM Transactions on Human–Robot Interaction (THRI), 11(2): 13:1–13:??, June 2022. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3488242.

Soh:2015:LAD

[SD15] Harold Soh and Yiannis Demiris.

Learning assistance by demonstration: smart mobility with shared control and paired haptic controllers. Journal of Human–Robot Interaction (JHRI), 4(3):76–100, December 2015. ISSN 2163-0364.

                                        URL https://dl.acm.org/
                                        doi/abs/10.5898/JHRI.4.3.
                                        Soh.

Short:2017:UAI

[SDFSM17] Elaine Schaertl Short, Eric C. Deng, David Feil-Seifer, and Maja J. Matarić. Understanding agency in interactions between children with autism and socially assistive robots. Journal of Human–Robot Interaction (JHRI), 6(3):21–47, December 2017. ISSN 2163-

                                        0364. URL https://dl.acm.
                                        org/doi/abs/10.5898/JHRI.
                                        6.3.Short.

Setter:2015:HIM

[SFEK15] Tina Setter, Alex Fouraker, Magnus Egerstedt, and Hiroaki Kawashima. Haptic interactions with multi-robot swarms using manipulability. Journal of Human–Robot Interaction (JHRI), 4(1):60–74, July 2015. ISSN 2163-0364.

                                        URL https://dl.acm.org/
                                        doi/abs/10.5898/JHRI.4.1.
                                        Setter.

Schrum:2023:CTL

[SGHb^{+}23] Mariah Schrum, Muyleng Ghuy, Erin Hedlund-botti, Manisha Natarajan, Michael Johnson, and Matthew Gombolay. Concerning trends in Likert scale usage in human-robot interaction: Towards improving best practices. ACM Transactions on Human–Robot Interaction (THRI), 12(3):33:1–33:??, September 2023. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3572784.

Shi:2022:TPA

[SGJ^{+}22] Zhonghao Shi, Thomas R.

Groechel, Shomik Jain, Kourtney Chima, Ognjen (Oggi) Rudovic, and Maja J. Matarić. Toward personalized affectaware socially assistive robot tutors for long-term interventions with children with autism. ACM Transactions on Human– Robot Interaction (THRI), 11 (4):39:1–39:??, December 2022. CODEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3526111.

Suguitan:2019:BHO

[SH19] Michael Suguitan and Guy Hoffman. Blossom: a handcrafted open-source robot. ACM Transactions on Human– Robot Interaction (THRI), 8(1): 2:1–2:27, March 2019. CO-DEN ???? ISSN 2573-9522 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/
3310356.

Spatola:2021:CIA

[SH21] Nicolas Spatola and Pascal Huguet. Cognitive impact of anthropomorphized robot gaze: Anthropomorphic gaze as social cues. ACM Transactions on Human–Robot Interaction (THRI), 10(4):35:1– 35:14, December 2021. CODEN ???? ISSN 2573-9522 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3459994.

Schodde:2020:AEE

[SHSK20] Thorsten Schodde, Laura Hoffmann, Sonja Stange, and Stefan Kopp. Adapt, explain, engage — a study on how social robots can scaffold second-language learning of children. ACM Transactions on Human–Robot Interaction (THRI), 9(1):6:1–6:27, January 2020. CODEN ???? ISSN 2573-9522 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3366422.

Sabanovic:2018:ATH

[SJ18] Selma Sabanović and Odest Chadwicke Jenkins. ACM Transactions on Human–Robot Interaction: a welcome from the Editors-in-Chief. ACM Transactions on Human–Robot Interaction (THRI), 7(1):1e:1– 1e:2, May 2018. CODEN ???? ISSN 2573-9522 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3209977.

Sabanovic:2019:E

[SJ19a] Selma Sabanović and Odest Chadwicke Jenkins. Editorial. ACM Transactions on Human–Robot Interaction (THRI), 8(3):13:1, August 2019. CODEN ???? ISSN 2573-9522 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3345004.

Sabanovic:2019:EHS

[SJ19b] Selma Sabanović and Odest Chadwicke Jenkins. Editorial — the HRI spring: Shaping our future from a foundation of diverse scholarship. ACM Transactions on Human–Robot Interaction (THRI), 8(1):1:1–1:2, March 2019. CODEN ???? ISSN 2573-9522 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3317314.

Sabanovic:2022:ATH

[SJ22] Selma Sabanović and Odest Chadwicke Jenkins. ACM Transactions on Human-Robot Interaction: The state of the journal. ACM Transactions on Human–Robot Interaction (THRI), 11(1):1e:1–1e:3, March 2022. CODEN ???? ISSN 2573-9522 (electronic). URL

https://dl.acm.org/doi/10.
1145/3488567.

Strabala:2013:TSH

[SLD^{+}13] Kyle Strabala, Min Kyung Lee, Anca Dragan, Jodi Forlizzi, Siddhartha S. Srinivasa, Maya Cakmak, and Vincenzo Micelli. Toward seamless human–robot handovers. Journal of Human–Robot Interaction (JHRI), 2(1):112–132, February 2013. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
2.1.Strabala.

Schneider:2022:SIM

[SLTK22] Sebastian Schneider, Yuyi Liu, Kanako Tomita, and Takayuki Kanda. Stop ignoring me! On fighting the trivialization of social robots in public spaces. ACM Transactions on Human– Robot Interaction (THRI), 11 (2):11:1–11:??, June 2022. CO-DEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3488241.

St-Onge:2019:ERS

[SOCAG^{+}19] David St-Onge, Ulysse Côté-Allard, Kyrre Glette, Benoit Gosselin, and Giovanni Beltrame. Engaging with robotic swarms: Commands from expressive motion. ACM Transactions on Human–Robot Interaction (THRI), 8(2):11:1– 11:26, June 2019. CODEN ???? ISSN 2573-9522 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3323213.

Spitale:2022:SAR

[SOG^{+}22] Micol Spitale, Sarah Okamoto, Mahima Gupta, Hao Xi, and Maja J. Matarić. Socially assistive robots as storytellers that elicit empathy. ACM Transactions on Human–Robot Interaction (THRI), 11(4):46:1– 46:??, December 2022. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3538409.

Salvini:2021:SMR

[SPGB21] Pericle Salvini, Diego Paez-Granados, and Aude Billard. On the safety of mobile robots serving in public spaces: Identifying gaps in EN ISO 13482:2014 and calling for a new standard. ACM Transactions on Human–Robot Interaction (THRI), 10(3):19:1–19:27, July 2021. CODEN ???? ISSN 2573-9522 (electronic). URL

https://dl.acm.org/doi/10.
1145/3442678.

Schadenberg:2021:SWY

[SRHE21] Bob R. Schadenberg, Dennis Reidsma, Dirk K. J. Heylen, and Vanessa Evers. “I see what you did there”: Understanding people’s social perception of a robot and its predictability. ACM Transactions on Human–Robot Interaction (THRI), 10(3):28:1– 28:28, July 2021. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3461534.

Sabanovic:2014:DRW

[SRK14] Selma Sabanović, Sarah M.

Reeder, and Bobak Kechavarzi. Designing robots in the wild: in situ prototype evaluation for a break management robot. Journal of Human–Robot Interaction (JHRI), 3(1):70–88, February 2014. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
3.1.Sabanovic.

Sandini:2018:HRR

[SS18] Giulio Sandini and Alessandra Sciutti. Humane robotsfrom robots with a humanoid body to robots with an anthropomorphic mind. ACM Transactions on Human–Robot Interaction (THRI), 7(1):7:1– 7:4, May 2018. CODEN ???? ISSN 2573-9522 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3208954.

Stefanov:2019:MHV

[SSK^{+}19] Kalin Stefanov, Giampiero Salvi, Dimosthenis Kontogiorgos, Hedvig Kjellström, and Jonas Beskow. Modeling of human visual attention in multiparty open-world dialogues. ACM Transactions on Human– Robot Interaction (THRI), 8 (2):8:1–8:21, June 2019. CO-DEN ???? ISSN 2573-9522 (electronic). URL https://dl.

                                        acm.org/doi/abs/10.1145/
                                        3323231.

Salomons:2021:MOA

[SSQS21] Nicole Salomons, Sarah Strohkorb Sebo, Meiying Qin, and Brian Scassellati. A minority of one against a majority of robots: Robots cause normative and informational conformity. ACM Transactions on Human–Robot Interaction (THRI), 10(2):15:1–15:22, May 2021. CODEN ???? ISSN 2573-9522 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3442627.

Sakamoto:2021:IMI

[SST21] Takafumi Sakamoto, Akihito Sudo, and Yugo Takeuchi. Investigation of model for initial phase of communication: Analysis of humans interaction by robot. ACM Transactions on Human–Robot Interaction (THRI), 10(2):14:1– 14:27, May 2021. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3439719.

Schulz:2019:ATH

[STH19] Trenton Schulz, Jim Torresen, and Jo Herstad. Animation techniques in human–robot interaction user studies: a systematic literature review. ACM Transactions on Human–Robot Interaction (THRI), 8(2):12:1– 12:22, June 2019. CODEN ???? ISSN 2573-9522 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3317325.

Sharma:2022:AHL

[SV22] Medha Sharma and Kavita Vemuri. Accepting human-like avatars in social and professional roles. ACM Transactions on Human–Robot Interaction (THRI), 11(3):28:1–28:??, September 2022. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3526026.

Seibt:2021:CHS

[SVD21] Johanna Seibt, Christina Vestergaard, and Malene F. Damholdt. The complexity of human social interactions calls for mixed methods in HRI: Comment on “A Primer for Conducting Experiments in Human–robot Interaction,” by G. Hoffman and X. Zhao. ACM Transactions on Human–Robot Interaction (THRI), 10(1):10:1– 10:4, February 2021. CO-DEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/

3439715. See [HZ21].

Seifi:2023:FHI

[SVKF23] Hasti Seifi, Steven A. Vasquez, Hyunyoung Kim, and Pooyan Fazli. First-hand impressions: Charting and predicting user impressions of robot hands. ACM Transactions on Human– Robot Interaction (THRI), 12 (3):35:1–35:??, September 2023. CODEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3580592.

Siebinga:2022:HFA

[SZA22] Olger Siebinga, Arkady Zgonnikov, and David Abbink. A human factors approach to validating driver models for interactionaware automated vehicles. ACM Transactions on Human– Robot Interaction (THRI), 11 (4):47:1–47:??, December 2022. CODEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3538705.

Tanaka:2015:RAR

[Tan15] Yoshiyuki Tanaka. Robot-aided rehabilitation methodology for enhancing movement smoothness by using a human hand trajectory generation model with task-related constraints. Journal of Human–Robot Interaction (JHRI), 4(3):101–119, December 2015. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
4.3.Tanaka.

Taylor:2020:RCP

[TCR20] Angelique Taylor, Darren M.

Chan, and Laurel D. Riek. Robot-centric perception of human groups. ACM Transactions on Human–Robot Interaction (THRI), 9(3):15:1– 15:21, July 2020. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3375798.

Thellman:2022:MSA

[TdGZ22] Sam Thellman, Maartje de Graaf, and Tom Ziemke. Mental state attribution to robots: a systematic review of conceptions, methods, and findings. ACM Transactions on Human– Robot Interaction (THRI), 11 (4):41:1–41:??, December 2022. CODEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3526112.

Takayama:2012:III

[TGS12] Leila Takayama, Willow Garage, and Brian Scassellati. Introduction to the inaugural issue: a special issue on interdisciplinary work at the intersection of systems and human sciences. Journal of Human– Robot Interaction (JHRI), 1(1): 2–3, July 2012. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI/
1.1.Takayama.

Trafton:2013:ARE

[THH^{+}13] J. Gregory Trafton, Laura M. Hiatt, Anthony M. Harrison, Franklin P. Tamborello, Sangeet S. Khemlani, and Alan C. Schultz. ACT-R/E: an embodied cognitive architecture for human–robot interaction. Journal of Human–Robot Interaction (JHRI), 2(1):30–55, February 2013. ISSN 2163-

                                        0364. URL https://dl.acm.
                                        org/doi/abs/10.5898/JHRI.
                                        2.1.Trafton.

Toris:2014:RMS

[TKC14] Russell Toris, David Kent, and Sonia Chernova. The robot management system: a framework for conducting human–robot interaction studies through crowdsourcing. Journal of Human–Robot Interaction (JHRI), 3(2):25–49, July 2014. ISSN 2163-0364.

                                        URL https://dl.acm.org/
                                        doi/abs/10.5555/3109829.
                                        3109831.

Tanaka:2012:CTC

[TM12] Fumihide Tanaka and Shizuko Matsuzoe. Children teach a care-receiving robot to promote their learning: field experiments in a classroom for vocabulary learning. Journal of Human–Robot Interaction (JHRI), 1(1):78–95, July 2012. ISSN 2163-0364.

                                        URL https://dl.acm.org/
                                        doi/abs/10.5898/JHRI.1.1.
                                        Tanaka.

Tian:2021:TSE

[TO21] Leimin Tian and Sharon Oviatt. A taxonomy of social errors in human–robot interaction. ACM Transactions on Human–Robot Interaction (THRI), 10(2):13:1–13:32, May 2021. CODEN ???? ISSN 2573-9522 (electronic). URL

https://dl.acm.org/doi/10.
1145/3439720.

Torta:2013:ATS

[TOW^{+}13] Elena Torta, Johannes Oberzaucher, Franz Werner, Raymond H. Cuijpers, and James F. Juola. Attitudes towards socially assistive robots in intelligent homes: results from laboratory studies and field trials. Journal of Human–Robot Interaction (JHRI), 1(2):76–99, February 2013. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
1.2.Torta.

Trafton:2021:PT

[TRK21] J. Gregory Trafton, Paula Raymond, and Sangeet Khemlani. The power of theory. ACM Transactions on Human– Robot Interaction (THRI), 10 (1):11:1–11:3, February 2021. CODEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3439716.

Thellman:2021:PBP

[TZ21] Sam Thellman and Tom Ziemke. The perceptual belief problem: Why explainability is a tough challenge in social robotics. ACM Transactions on Human–Robot Interaction (THRI), 10(3):29:1–29:15, July 2021. CODEN ???? ISSN 2573-9522 (electronic). URL

https://dl.acm.org/doi/10.
1145/3461781.

Urakami:2023:NCH

[US23] Jacqueline Urakami and Katie Seaborn. Nonverbal cues in human-robot interaction: a communication studies perspective. ACM Transactions on Human–Robot Interaction (THRI), 12(2):22:1–22:??, June 2023. CODEN ???? ISSN 2573-9522 (electronic). URL

https://dl.acm.org/doi/10.
1145/3570169.

vandenBrule:2016:WSP

[vdBBD^{+}16] Rik van den Brule, Gijsbert Bijlstra, Ron Dotsch, Pim Haselager, and Daniël H. J. Wigboldus. Warning signals for poor performance improve human–robot interaction. Journal of Human–Robot Interaction (JHRI), 5(2):69–89, September 2016. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
5.2.Van_den_Brule.

Veloso:2018:IFO

[Vel18] Manuela M. Veloso. The increasingly fascinating opportunity for human–robot-AI interaction: The CoBot mobile service robots. ACM Transactions on Human–Robot Interaction (THRI), 7(1):5:1–5:2, May 2018. CODEN ???? ISSN 2573-9522 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3209541.

Venture:2019:REM

[VK19] Gentiane Venture and Dana Kulić. Robot expressive motions: a survey of generation and evaluation methods. ACM Transactions on Human–Robot Interaction (THRI), 8(4):20:1– 20:17, December 2019. CO-DEN ???? ISSN 2573-9522 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/
3344286.

Vignolo:2022:HRE

[VPR^{+}22] Alessia Vignolo, Henry Powell, Francesco Rea, Alessandra Sciutti, Luke Mcellin, and John Michael. A humanoid robot’s effortful adaptation boosts partners’ commitment to an interactive teaching task. ACM Transactions on Human–Robot Interaction (THRI), 11(1):9:1–9:17, March 2022. CODEN ???? ISSN 2573-9522 (electronic). URL

https://dl.acm.org/doi/10.
1145/3481586.

vanStraten:2020:TAR

[vSPKB20] Caroline L. van Straten, Jochen Peter, Rinaldo Kühne, and Alex Barco. Transparency about a robot’s lack of human psychological capacities: Effects on child–robot perception and relationship formation. ACM Transactions on Human–Robot Interaction (THRI), 9(2):11:1– 11:22, February 2020. CO-DEN ???? ISSN 2573-9522 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/
3365668.

Valner:2022:UMR

[VWKP22] Robert Valner, Selma Wanna, Karl Kruusamäe, and Mitch Pryor. Unified meaning representation format (UMRF)

— a task description and execution formalism for HRI. ACM Transactions on Human– Robot Interaction (THRI), 11 (4):38:1–38:??, December 2022. CODEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3522580.

Washburn:2020:REP

[WAAR20] Auriel Washburn, Akanimoh Adeleye, Thomas An, and Laurel D. Riek. Robot errors in proximate HRI: How functionality framing affects perceived reliability and trust. ACM Transactions on Human–Robot Interaction (THRI), 9(3):19:1– 19:21, July 2020. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3380783.

Wagner:2015:RSC

[Wag15] Alan R. Wagner. Robots that stereotype: creating and using categories of people for human– robot interaction. Journal of Human–Robot Interaction (JHRI), 4(2):97–127, September 2015. ISSN 2163-0364.

URL https://dl.acm.org/
doi/abs/10.5555/3109842.
3109847.

Williams:2015:CRR

[WBS15] Tom Williams, Priscilla Briggs, and Matthias Scheutz. Covert robot–robot communication: human perceptions and implications for human–robot interaction. Journal of Human– Robot Interaction (JHRI), 4(2): 24–49, September 2015. ISSN 2163-0364. URL https://dl.

acm.org/doi/abs/10.5898/
JHRI.4.2.Williams.

Wareham:2023:SCD

[WdHVvR23] Todd Wareham, Ronald de Haan, Andrew Vardy, and Iris van Rooij. Swarm control for distributed construction: a computational complexity perspective. ACM Transactions on Human–Robot Interaction (THRI), 12(1):6:1–6:??, March 2023. CODEN ???? ISSN 2573-9522 (electronic). URL

https://dl.acm.org/doi/10.
1145/3555078.

Wijnen:2020:NWT

[WDR^{+}20] Frances M. Wijnen, Daniel P.

Davison, Dennis Reidsma, Jan Van Der Meij, Vicky Charisi, and Vanessa Evers. Now we’re talking: Learning by explaining your reasoning to a social robot. ACM Transactions on Human–Robot Interaction (THRI), 9(1):5:1–5:29, January 2020. CODEN ???? ISSN 2573-9522 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3345508.

Wullenkord:2020:IRN

[WE20] Ricarda Wullenkord and Friederike Eyssel. The influence of robot number on robot group perception — a call for action. ACM Transactions on Human–Robot Interaction (THRI), 9(4):27:1– 27:14, October 2020. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3394899.

Wrede:2013:USK

[WEG^{+}13] Sebastian Wrede, Christian Emmerich, Ricarda Grünberg, Arne Nordmann, Agnes Swadzba, and Jochen Steil. A user study on kinesthetic teaching of redundant robots in task and configuration space. Journal of Human–Robot Interaction (JHRI), 2(1):56–81, February 2013. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
2.1.Wrede.

Wen:2023:CNB

[WKP^{+}23] Ruchen Wen, Boyoung Kim, Elizabeth Phillips, Qin Zhu, and Tom Williams. Comparing norm-based and rolebased strategies for robot communication of role-grounded moral norms. ACM Transactions on Human–Robot Interaction (THRI), 12(3):30:1–30:??, September 2023. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3571719.

Winkle:2023:YWM

[WLTO23] Katie Winkle, Erik Lagerstedt, Ilaria Torre, and Anna Offenwanger. 15 years of (who)man robot interaction: Reviewing the H in humanrobot interaction. ACM Transactions on Human–Robot Interaction (THRI), 12(3):28:1– 28:??, September 2023. CO-DEN ???? ISSN 2573-9522 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3571718.

Wong:2016:AAD

[WM16] Richmond Y. Wong and Deirdre K.

Mulligan. These aren’t the autonomous drones you’re looking for: investigating privacy concerns through concept videos. Journal of Human–Robot Interaction (JHRI), 5(3):26–54, December 2016. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
5.3.Wong.

Walker:2015:UCD

[WML15] Amber M. Walker, David P.

Miller, and Chen Ling. Usercentered design of an attitudeaware controller for ground reconnaissance robots. Journal of Human–Robot Interaction (JHRI), 4(1):30–59, July 2015. ISSN 2163-0364.

URL https://dl.acm.org/
doi/abs/10.5898/JHRI.4.1.
Walker.

Walker:2023:VAM

[WPC^{+}23] Michael Walker, Thao Phung, Tathagata Chakraborti, Tom Williams, and Daniel Szafir. Virtual, augmented, and mixed reality for human-robot interaction: a survey and virtual design element taxonomy. ACM Transactions on Human– Robot Interaction (THRI), 12 (4):43:1–43:??, 2023. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3597623.

Whittaker:2021:DPE

[WRPZ21] Steve Whittaker, Yvonne Rogers, Elena Petrovskaya, and Hongbin Zhuang. Designing personas for expressive robots: Personality in the new breed of moving, speaking, and colorful social home robots. ACM Transactions on Human–Robot Interaction (THRI), 10(1):8:1– 8:25, February 2021. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3424153.

Westerberg:2013:VEB

[WS13] Simon Westerberg and Anton Shiriaev. Virtual environmentbased teleoperation of forestry machines: designing future interaction methods. Journal of Human–Robot Interaction (JHRI), 2(3):84–110, September 2013. ISSN 2163-0364.

URL https://dl.acm.org/
doi/abs/10.5898/JHRI.2.3.
Westerberg.

Wallbridge:2022:EDP

[WSG^{+}22] Christopher D. Wallbridge, Alex Smith, Manuel Giuliani, Chris Melhuish, Tony Belpaeme, and Séverin Lemaignan. The effectiveness of dynamically processed incremental descriptions in human robot interaction. ACM Transactions on Human–Robot Interaction (THRI), 11(1):7:1–7:24, March 2022. CODEN ???? ISSN 2573-9522 (electronic). URL

https://dl.acm.org/doi/10.
1145/3481628.

Wallkotter:2021:EEA

[WTC^{+}21] Sebastian Wallkötter, Silvia Tulli, Ginevra Castellano, Ana Paiva, and Mohamed Chetouani. Explainable embodied agents through social cues: a review. ACM Transactions on Human–Robot Interaction (THRI), 10(3):27:1–27:24, July 2021. CODEN ???? ISSN 2573-9522 (electronic). URL

https://dl.acm.org/doi/10.
1145/3457188.

Wilson:2020:CDF

[WTDS20] Jason R. Wilson, Linda Tickle-Degnen, and Matthias Scheutz. Challenges in designing a fully autonomous socially assistive robot for people with Parkinson’s disease. ACM Transactions on Human–Robot Interaction (THRI), 9(3):20:1– 20:31, July 2020. CODEN ???? ISSN 2573-9522 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3379179.

Williams:2023:CMC

[WYJ23] Katherine J. Williams, Madeleine S.

Yuh, and Neera Jain. A computational model of coupled human trust and self-confidence dynamics. ACM Transactions on Human–Robot Interaction (THRI), 12(3):39:1–39:??, September 2023. CODEN ???? ISSN 2573-9522 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3594715.

Young:2017:HGC

[You17] James Everett Young. An HRI graduate course for exposing technologists to the importance of considering social aspects of technology. Journal of Human–Robot Interaction (JHRI), 6(2):27–47, September 2017. ISSN 2163-0364.

                                        URL https://dl.acm.org/
                                        doi/abs/10.5898/JHRI.6.2.
                                        Young.

Yogeeswaran:2016:IER

[YZL^{+}16] Kumar Yogeeswaran, Jakub Złotowski, Megan Livingstone, Christoph Bartneck, Hidenobu Sumioka, and Hiroshi Ishiguro. The interactive effects of robot anthropomorphism and robot ability on perceived threat and support for robotics research. Journal of Human–Robot Interaction (JHRI), 5(2):29–47, September 2016. ISSN 2163-

                                        0364. URL https://dl.acm.
                                        org/doi/abs/10.5898/JHRI.
                                        5.2.Yogeeswaran.

Zhou:2020:MPS

[ZCG^{+}20] Tian Zhou, Jackie S. Cha, Glebys Gonzalez, Juan P. Wachs, Chandru P. Sundaram, and Denny Yu. Multimodal physiological signals for workload prediction in robot-assisted surgery. ACM Transactions on Human–Robot Interaction (THRI), 9(2):12:1–12:26, February 2020. CODEN ???? ISSN 2573-9522 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3368589.

Zenk:2017:USR

[ZCV^{+}17] J. D. Zenk, C. R. Crowell, M. Villano, J. Kaboski, K. Tang, and J. Diehl. Unconventional students in robotics and HRI education: a review of two initiatives. Journal of Human–Robot Interaction (JHRI), 6(2):92–110, September 2017. ISSN 2163-

0364. URL https://dl.acm.
org/doi/abs/10.5898/JHRI.
6.2.Zenk.

Zhou:2016:CST

[ZCW^{+}16] Tian Zhou, Maria E. Cabrera, Juan P. Wachs, Thomas Low, and Chandru Sundaram. A comparative study for telerobotic surgery using free hand gestures. Journal of Human– Robot Interaction (JHRI), 5(2): 1–28, September 2016. ISSN 2163-0364. URL https://dl.

acm.org/doi/abs/10.5898/
JHRI.5.2.Zhou.

Zhang:2020:EAU

[ZFS^{+}20] Wenjuan Zhang, David Feltner, James Shirley, David Kaber, and Manida S. Neubert. Enhancement and application of a UAV control interface evaluation technique: Modified GEDIS-UAV. ACM Transactions on Human–Robot Interaction (THRI), 9(2):14:1–14:20, February 2020. CODEN ???? ISSN 2573-9522 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3368943.

Zhou:2019:IDL

[ZZKC19] Jin Zhou, Haibei Zhu, Minwoo Kim, and Mary L. Cummings. The impact of different levels of autonomy and training on operators’ drone control strategies. ACM Transactions on Human–Robot Interaction (THRI), 8(4):22:1–22:15, December 2019. CODEN ???? ISSN 2573-9522 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3344276.